### v1

# MRAG-Suite: A Diagnostic Evaluation Platform for Visual Retrieval-Augmented Generation

Abstract
Multimodal Retrieval-Augmented Generation (Visual RAG) significantly advances question answering by integrating visual and textual evidence. Yet, current evaluations fail to systematically account for query difficulty and ambiguity. We propose MRAG-Suite, a diagnostic evaluation platform integrating diverse multimodal benchmarks (WebQA, Chart-RAG, Visual-RAG, MRAG-Bench). We introduce difficulty-based and ambiguity-aware filtering strategies, alongside MM-RAGChecker, a claim-level diagnostic tool. Our results demonstrate substantial accuracy reductions under difficult and ambiguous queries, highlighting prevalent hallucinations. MM-RAGChecker effectively diagnoses these issues, guiding future improvements in Visual RAG systems.
MRAG-Suite: A Diagnostic Evaluation Platform for Visual Retrieval-Augmented Generation
Yuelyu Ji University of Pittsburgh yuj49@pitt.edu
1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a foundational paradigm for enhancing the factuality and reliability of large language models (LLMs) by grounding their outputs in external evidence. With recent advancements in multimodal capabilities, visual RAG systems—capable of leveraging textual and visual information simultaneously have garnered significant interest, promising substantial improvements in answering complex questions grounded in diverse modalities such as charts, plots, figures, images, and documents. Despite these advancements, current evaluation methodologies have not kept pace with these multimodal developments, primarily focusing either on purely textual contexts or closed-domain multimodal setups that provide pre-selected visual evidence. Consequently, the research community lacks rigorous evaluation benchmarks capable of systematically assessing multimodal retrieval accuracy, reasoning fidelity, and robustness to ambiguity and distractors.
Figure 1: Illustration of our multimodal claim verification process. Given a long-form answer generated by a retrieval-augmented generation (RAG) system, we extract atomic factual claims and assess whether each is supported by retrieved visual evidence. Claims are categorized as Entailed if supported by at least one image. In this example, the model correctly describes several physical traits and habitat features of a marmot , but omits the gold claim about “high-altitude environments” , which remains unsupported by the retrieved evidence.
Existing multimodal benchmarks like ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and MMMU-Pro (Yue et al., 2024) typically provide necessary visual contexts as part of the inputs, neglecting the retrieval component critical in real-world scenarios. Meanwhile, retrieval-centric evaluations like Visual Haystacks (Wu et al., 2024) emphasize challenging retrieval tasks but limit their scope to a single domain without fine-grained error diagnostics. Recent text-focused diagnostic frameworks such as RAGChecker and RefChekcer (Ru et al., 2024; Hu et al., 2024b) offer detailed claim-level analysis but remain restricted to textual contexts, unable to handle the additional complexity introduced by visual evidence.
To address these critical gaps, we propose MRAG-Suite, a comprehensive diagnostic evaluation platform designed explicitly for visual RAG systems. MRAG-Suite integrates eight heterogeneous datasets, spanning a wide range of real-world modalities, including charts, plots, scanned documents, slides, and open-web images. Our benchmark systematically introduces difficulty-based filtering and ambiguity-aware query variants, coupled with controlled distractor scenarios, enabling robust assessment of a model’s capability to handle ambiguity, retrieve precise evidence, and produce faithful, evidence-grounded answers.
Specifically, MRAG-Suite contributes:
  1. 1.
    Multi-domain multimodal benchmark: We consolidate and normalize datasets across diverse domains, providing unified evaluation standards and enabling direct comparisons of model performance across various multimodal tasks.
  2. 2.
    Difficulty- and ambiguity-aware filtering: By filtering trivial or easily guessable questions and introducing ambiguity through carefully constructed query variations, MRAG-Suite rigorously evaluates robustness against ambiguity and query complexity.
  3. 3.
    Claim-level multimodal diagnostic tool (MM-RAGChecker): Extending claim-level diagnostics to multimodal contexts, MM-RAGChecker systematically verifies each claim within generated answers against textual and visual evidence, enabling precise identification of hallucinations, modality biases, and retrieval accuracy.
All code, prompts, and diagnostics are available at:
https://anonymous.4open.science/status/MRAGChecker-B33D
2 Related Work
Retrieval-Augmented Generation Evaluation
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (Ru et al., 2024) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
Multimodal QA and Reasoning Benchmarks
Benchmarks such as ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and their advanced variants MMMU-Pro (Yue et al., 2024) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu et al., 2024) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsLi et al. (2025b, a).
Biases in Multimodal Retrieval-Augmented Systems
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao et al., 2025) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlyWang et al. (2025); Liang et al. (2025); Jiang et al. (2024); Dan et al. (2024); Lu et al. (2025); Li et al. .
Domain-specific Multimodal QA
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (Yang et al., 2025) and VDocRAG (Tanaka et al., 2025), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
3 MRAG-Suite Benchmark Construction
3.1 Task Definition
We evaluate multimodal Retrieval-Augmented Generation (RAG) for open-ended QA. Given a query $q$, the system retrieves textual ($E^{\text{txt}}$) and visual ($E^{\text{img}}$) evidence and must produce (i) a concise short answer and (ii) a grounded long answer: $(a^{\text{short}},a^{\text{long}})=f\!\big(q,E^{\text{txt}},E^{\text{img}}\big).$
3.2 Corpora Overview
MRAG-Suite merges eight heterogeneous sources spanning charts, natural photos, open-web images, scanned PDFs, and slides. Each source contributes distinct modalities and reasoning demands:
  • WebQA Chang et al. (2022): open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG Yang et al. (2025)(from ChartQA Masry et al. (2022)): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu et al. (2025): queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench Hu et al. (2024a): 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu et al. (2024) subsets: Arxiv Li et al. (2024) (scholarly PDF pages), Plot Methani et al. (2019) (technical plots), Slide Tanaka et al. (2023) (slide decks), and Document Tito et al. (2022) (scanned manuals/forms). Each page is treated as a visual document with OCR text.
Dataset statistics.
Table 1 summarizes domains, filtering counts, image resolutions, and text length statistics.
Table 1: Overview of the eight sources in MRAG-Suite, covering domains, filtering counts, image resolution, and question/answer lengths. The suite spans open-web images, charts/plots, scanned documents, slides, and scholarly PDFs to enable cross-domain visual RAG evaluation.
Dataset Domain Two-way filter Avg. Img (px) Max Q Max A Avg Q Avg A Ambig.#
[…]
Figure 2: $M$-RAGChecker overview: claim extraction, multimodal evidence matching, and verification with per-claim labels and diagnostic metrics.
3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captions; text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.
3.4 Two-Step Filtering for Non-triviality
We remove trivial and guessable items via:
  1. 1.
    Retrieval-Independent Filter: questions solved by a strong closed-book model (Claude 3.5, confidence >0.9>0.9) or whose answers appear verbatim in the question.
  2. 2.
    Difficulty-Based Filter: rank remaining items by multi-hop requirement, modality dependency, and baseline success rate; drop the easiest $\sim$10% per domain.
We release two query variants: Filtered (minimal, disambiguated) and Full (original, potentially noisy).
Ambiguity-aware evaluation.
To isolate the impact of query ambiguity on Visual RAG, we construct an ambiguity-aware subset (200 items) using a two-stage pipeline: an LLM-based pre-filter with short rationales, followed by double-annotator adjudication (substantial agreement, $\kappa\!=\!0.74$). We also provide evidence-grounded answer rewriting and disambiguated query rewrites for these items. Full prompts and guidelines are deferred to Appendix A.1.
3.5 Result of Distractors
To probe robustness, we evaluate three settings:
  • gt_only (GO): only gold evidence;
  • gt_plus_distractors (GPD): gold mixed with CLIP-similar distractors;
  • distractors_only (DO): only distractors, testing spurious reasoning.
3.6 Output Specification
Systems must return both short and long answers. We compute EM/Accuracy for $a^{\text{short}}$, and claim-level diagnostics plus ROUGE-L for $a^{\text{long}}$ (MM-RAGChecker in Sec. 4).
Having standardized data and retrieval settings, we next introduce MM-RAGChecker, a claim-level diagnostic tool that evaluates how long-form answers are grounded in the retrieved multimodal evidence. It operates in three stages—claim extraction, evidence matching, and verification—and produces a set of fine-grained metrics for hallucination, faithfulness, and cross-modality behavior.
Evaluation axes (metric families).
We evaluate along two complementary metric families rather than by answer length: (i) End-task accuracy on the short answer ($a^{\text{short}}$) via EM/Acc; and (ii) Claim-level diagnostics on the long answer ($a^{\text{long}}$) using MM-RAGChecker ; Hallucination, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, and cross-modality metrics).
4 MM-RAGChecker: Multimodal Claim-Level Diagnosis
4.1 Overview
Stage 1 — Claim extraction. We split the long answer $a^{\text{long}}$ into minimal verifiable units $\mathcal{C}=\{c_{i}\}$ using simple heuristics plus an LLM splitter for compound clauses (e.g., numeric comparisons).
Stage 2 — Per-evidence judging (3-way). Given images $\{I_{k}\}_{k=1}^{K_{\mathrm{img}}}$ and text passages $\{t_{k}\}_{k=1}^{K_{\mathrm{txt}}}$ retrieved for the question, we directly judge each pair $(c_{i},e)$, $e\in\{I_{k}\}\cup\{t_{k}\}$, with a three-way decision $\{\textsc{Entailment},\textsc{Neutral},\textsc{Contradiction}\}$. No image captions are synthesized and no numeric scores are used. We cap to top $K_{\mathrm{img}}{=}3$ and $K_{\mathrm{txt}}{=}3$ by default and set temperature to zero for determinism. For each claim we store two multisets of judgments $\mathcal{J}^{\mathrm{img}}_{i}=\{j(c_{i},I_{k})\}$ and $\mathcal{J}^{\mathrm{txt}}_{i}=\{j(c_{i},t_{k})\}$.
Stage 3 — Aggregation to claim labels. Each claim receives a single final label $L_{i}\in\{\textsc{Entailment},\textsc{Neutral},\textsc{Contradiction}\}$ via a simple precedence rule: (i) $L_{i}{=}\textsc{Entailment}$ if any $j\in\mathcal{J}^{\mathrm{img}}_{i}\cup\mathcal{J}^{\mathrm{txt}}_{i}$ is Entailment; (ii) else $L_{i}{=}\textsc{Contradiction}$ if any judgment is Contradiction; (iii) else $L_{i}{=}\textsc{Neutral}$. For modality analysis, define binary flags $s^{\mathrm{img}}_{i}{=}1$ iff some $j\in\mathcal{J}^{\mathrm{img}}_{i}$ is Entailment, and analogously $s^{\mathrm{txt}}_{i}$ for text. We also mark a retrieved item $e$ as used if there exists a claim $c_{i}$ with $j(c_{i},e){=}\textsc{Entailment}$. All metrics below are computed on these final labels and flags.
4.2 Metrics
Core metrics.
Let $\mathcal{C}$ be the set of extracted claims, $L_{i}$ the final label of $c_{i}$, and $\mathcal{E}$ the set of retrieved items. Then
$\mathrm{Hallucination\ Rate}$ $=\frac{\#\{\,i:L_{i}=\textsc{Neutral}\,\}}{|\mathcal{C}|},$ (1)
$\mathrm{Faithfulness}$ $=\frac{\#\{\,i:L_{i}=\textsc{Entailment}\,\}}{|\mathcal{C}|}.$ (2)
Let $\mathcal{G}$ be gold salient claims mined from a reference long answer.
$\mathrm{Claim\ Recall\ (CR)}$ $=\frac{\#\{\text{gold }g\in\mathcal{G}\ \text{that appear in }a^{\text{long}}\}}{|\mathcal{G}|}.$ (3)
Let $\mathcal{E}_{\mathrm{used}}=\{e\in\mathcal{E}:\exists i,\ j(c_{i},e)=\textsc{Entailment}\}$.
$\mathrm{Context\ Precision\ (CP)}$ $=\tfrac{|\mathcal{E}_{\mathrm{used}}|}{|\mathcal{E}|}.$ (4)
Self-knowledge measures how many entailed claims did not rely on retrieval: $\mathrm{Self\text{-}Knowledge}=\frac{\#\{\textsc{Entailment }c_{i}\ \text{with no }j(c_{i},e)=\textsc{Entailment}\ \forall e\in\mathcal{E}\}}{\#\{\textsc{Entailment }c_{i}\}}$
Cross-modality metrics.
Let $K_{\mathrm{img}}$$K_{\mathrm{txt}}$ be the numbers of retrieved images/texts. Using $s^{\mathrm{img}}_{i},s^{\mathrm{txt}}_{i}\in\{0,1\}$:
$\Delta\mathrm{CR}$ $=\mathrm{CR}_{\mathrm{img}}-\mathrm{CR}_{\mathrm{txt}},$ (5)
$\Delta\mathrm{CP}$ $=\mathrm{CP}_{\mathrm{img}}-\mathrm{CP}_{\mathrm{txt}}.$ (6)
Coverage of each modality:
$\mathrm{VIS\!-\!Hit@}k$ $=\frac{\#\left\{\begin{subarray}{c}\text{retrieved images that}\\ \text{entail at least one claim}\end{subarray}\right\}}{K_{\mathrm{img}}},$ (7)
$\mathrm{TXT\!-\!MissRate}$ $=1-\frac{\#\left\{\begin{subarray}{c}\text{retrieved text chunks that}\\ \text{entail any claim}\end{subarray}\right\}}{K_{\mathrm{txt}}}.$ (8)
Agreement and attribution:
$\mathrm{CMA}$ $=\frac{\sum_{i}\mathbf{1}[s^{\mathrm{img}}_{i}=1\wedge s^{\mathrm{txt}}_{i}=1]}{\sum_{i}\mathbf{1}[s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1]},$ (9)
$\mathrm{V\!-\!HR}$ $=\frac{\#\{\,i:s^{\mathrm{img}}_{i}=1\wedge L_{i}=\textsc{Neutral}\,\}}{\#\{\,i:L_{i}=\textsc{Neutral}\,\}},$ (10)
$\mathrm{D\!-\!HR}$ $=\frac{\#\{\,i:s^{\mathrm{txt}}_{i}=1\wedge L_{i}=\textsc{Neutral}\,\}}{\#\{\,i:L_{i}=\textsc{Neutral}\,\}}.$ (11)
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 Pixtral-12B-2409 InternVL3-8B
Filt Full Filt Full Filt Full Filt Full Filt Full Filt Full
Chart-MRAG 22.4% 19.1% 15.3% 25.7% 10.8% 11.9% 11.2% 13.5% 14.7% 29.4% 19.2% 27.6%
MRAG-Bench 32.6% 33.8% 19.5% 24.6% 28.3% 26.1% 32.4% 30.7% 38.8% 37.5% 21.9% 21.4%
VisRAG-ArXiv 55.3% 48.7% 34.6% 34.5% 35.5% 35.1% 38.6% 36.9% 43.1% 43.7% 63.9% 55.4%
VisRAG-Doc 31.1% 35.4% 20.2% 21.9% 28.5% 29.2% 27.6% 28.2% 26.9% 26.6% 39.7% 34.6%
VisRAG-Plot 36.7% 45.8% 12.3% 13.6% 35.9% 36.8% 6.4% 9.2% 51.4% 56.5% 33.5% 36.1%
VisRAG-Slide 18.5% 25.3% 8.7% 10.4% 17.6% 22.1% 24.3% 22.6% 18.2% 15.5% 20.8% 22.9%
Visual-RAG 13.8% 13.1% 9.6% 11.5% 11.7% 10.6% 8.5% 9.2% 12.1% 13.3% 11.2% 12.6%
WebQA 16.2% 14.9% 12.7% 13.4% 14.8% 17.3% 20.6% 22.1% 26.4% 27.9% 17.8% 28.7%
Table 2: EM/Accuracy performance of each model on the Filtered (retained valid samples) and Full (all samples) splits.
Dataset Model Recall Precision F1 Halluc. Faith. Self‑know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
Chart-MRAG Qwen2.5-VL-72B 12.1/12.9 7.9/8.1 6.3/7.2 57.3/58.9 54.9/7.8 8.3/7.7 8.9/14.7 85.2/31.8
LLaVa-v1.6-34B 7.6/10.1 0.6/18.9 1.1/13.2 57.4/66.8 50.5/8.8 3.6/10.6 2.0/12.8 86.1/22.3
Claude 3.5 13.3/13.2 0.8/2.1 1.4/3.4 24.9/87.3 65.1/10.6 0.0/2.1 75/10 90/15
MRAG-Bench Qwen2.5-VL-72B 16.3 1.2 2.2 45.2 62.6 12.4 11.3 72.1
LLaVa-v1.6-34B 10.9 1.0 1.8 57.8 54.8 3.6 8.5 87.4
Claude 3.5 53.3 4.1 7.6 48.2 66.3 5.7 11.0 79.6
VisRAG-ArXiv Qwen2.5-VL-72B 23.7 1.9 3.5 30.3 67.2 17.2 23.2 61.0
LLaVa-v1.6-34B 19.7 1.7 3.1 37.0 65.5 12.3 19.4 67.4
Claude 3.5 17.0 1.6 2.9 41.4/0 64.4 9.1 16.9 71.6
VisRAG-Doc Qwen2.5-VL-72B 16.8 1.5 2.8 45.0 57.5 8.0 11.1 74.3
LLaVa-v1.6-34B 11.0 0.9 1.6 54.3 56.3 4.2 5.6 82.3
Claude 3.5 12.1 1.2 2.2 51.6 57.7 4.7 7.3 80.4
VisRAG-Plot Qwen2.5-VL-72B 17.7 1.9 3.4 42.1 65.5 9.6 16.4 75.1
LLaVa-v1.6-34B 9.5 0.8 1.5 60.9 50.1 3.2 1.9 88.6
Claude 3.5 17.1 1.5 2.8 46.7 62.8 8.3 13.5 75.9
VisRAG-Slide Qwen2.5-VL-72B 8.7 0.6 1.1 52.7 47.7 4.4 3.2 85.9
LLaVa-v1.6-34B 6.3 0.5 0.9 65.8 46.3 1.2 1.3 92.2
Claude 3.5 8.2 0.5 0.9 55.8 51.3 2.1 1.0 89.6
Visual-RAG Qwen2.5-VL-72B 6.5 0.5 1.0 74.0 42.3 2.7 5.2 11.0
LLaVa-v1.6-34B 5.3 0.4 0.8 76.8 44.1 3.2 4.5 12.0
Claude 3.5 6.0 0.6 1.1 85.1 39.5 1.8 6.1 14.5
WebQA Qwen2.5-VL-72B 8.8/8.1 0.8/16.1 1.5/10.8 62.8/70.4 52.5/7.9 6.5/3.6 5.9/13.5 89.7/30.4
LLaVa-v1.6-34B 6.5/7.1 0.7/15.8 1.3/9.8 64.5/70.8 50.2/6.2 2.5/4.1 1.1/12.2 91.4/28.6
Claude 3.5 9.0/7.9 1.1/1.8 1.7/3.7 26.6/98.8 62.7/10.7 0.3/0.6 72.0/8.8 87.4/13.6
Table 3: Diagnostic metrics on the Filtered split (image/text pairs where applicable).
Dataset Model $\DeltaCR Δ\DeltaCP$ VIS-Hit@k TXT-MissRate CMA V-HR D-HR
Chart-MRAG Phi-4 (GPD) 54.54 59.50 81.38 83.00 13.02 66.87 45.2
WebQA Phi-4 (GPD) 6.59 1.50 14.79 87.00 10.86 12.41 0.33
WebQA Claude 2.5 (GPD) 3.09 -0.35 13.00 83.33 6.52 32.44 3.70
[…]
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 InternVL3-8B Pixtral-12B-2409
DO GPD GO DO GPD GO DO GPD GO DO GPD GO DO GPD GO DO GPD GO
Chart-MRAG 22.7 25.4 26.8 10.3 18.1 13.9 8.2 15.5 9.7 15.9 19.8 17.4 23.3 19.5 21.7 14.8 14.1 16.6
MRAG-Bench 30.9 29.6 29.8 17.8 16.4 20.7 26.9 24.1 29.7 2.1 2.4 27.9 26.2 21.3 28.5 33.6 29.8 28.2
VisRAG-ArXiv 4.4 4.1 4.3 4.3 11.7 8.5 7.3 4.1 5.0 1.7 4.4 5.5 6.8 5.1 5.2 3.2 1.4 1.8
VisRAG-Doc 1.2 29.1 31.3 12.9 20.5 20.2 2.4 26.6 25.9 2.3 21.6 28.3 4.7 38.4 38.1 0.6 26.5 27.1
VisRAG-Plot 5.9 33.8 41.4 8.4 12.9 12.6 12.7 25.7 29.3 6.2 6.7 32.1 4.9 28.3 28.4 0.8 51.2 52.0
VisRAG-Slide 1.6 18.7 18.1 1.9 8.4 8.2 4.3 18.9 18.8 1.2 24.6 15.3 0.9 17.5 15.7 0.6 15.3 15.6
Visual-RAG 14.8 26.9 27.3 13.6 18.7 18.1 15.5 25.4 26.5 15.1 21.4 20.2 17.6 23.1 23.2 16.5 23.6 18.9
WebQA 17.2 15.8 15.4 12.3 12.6 13.4 9.9 14.3 16.4 17.5 7.3 12.8 20.1 20.8 17.4 25.6 25.3 0.9
Table 5: Scores (EM preferred; Accuracy used when EM is unavailable) under the three retrieval modes: DO = distractors_only, GPD = gt_plus_distractors, and GO = gt_only.
4.3 Retrieval Modes and Distractors
We compare distractors_only (DO), gt_plus_distractors (GPD), and gt_only (GO). Table 5 shows that DO degrades performance substantially; GPD lies between DO and GO. The GPD–GO gap quantifies susceptibility to distractors.
5 Experiments and Results
We evaluate MRAG-Suite along three axes: (1) end-task accuracy on short answers, (2) factuality and grounding quality of long answers via MM-RAGChecker, and (3) robustness to query difficulty, ambiguity, and distractors.
5.1 Experimental Setup
Models.
Our main baseline is LLaVa-v1.6-34B (vision-language model fine-tuned from LLaMA-2). Retrieval uses a CLIP-based dual encoder for images and a DPR retriever for text. At test time we fetch up to 5 passages and 5 images per query (when available) and serialize them (image features/captions + text) before generation. The model is prompted to output a short answer followed by a long answer. For comparison we also run Claude 3.5 (API) on a subset to estimate the upper bound of current proprietary VLMs.
Metrics.
For short answers we report EM/Accuracy. For long answers we compute ROUGE-L against our reference explanations and all claim-level diagnostics using MM-RAGChecker (Hallucination Rate, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, Modality Bias). We further report the fraction of retrieved items actually used (evidence utilization). Retrieval robustness is probed under three modes: gt_only (GO), gt_plus_distractors (GPD), and distractors_only (DO).
5.2 End-task Accuracy (short-answer EM/Acc)
Table 13 summarizes EM/Accuracy across datasets. On the Full split, LLaVA-2 reaches an overall accuracy of 62.5%. After applying our two-step filtering, accuracy drops to 47.8% (–14.7 points), confirming that the retained questions are substantially harder and more retrieval-dependent. Performance varies widely by domain: WebQA remains the easiest for the baseline (52% on the filtered set), while Chart-RAG is the hardest (low 30%s). Visual-RAG sits in between ( 45%), indicating persistent difficulty with fine-grained visual recognition. A no-retrieval ablation reduces accuracy to 25% on the filtered split, showing that at least half of the questions truly require external evidence. Manual inspection of 50 incorrect cases reveals two major categories: (a) Visual under-utilization (40%): the model retrieves the correct image/chart but misreads or ignores it; (b) Multi-hop failures (30%): the model retrieves partial evidence but fails to integrate it. The rest are simple reasoning mistakes or retriever misses.
5.3 Claim-level Diagnostics (MM-RAGChecker)
Using MM-RAGChecker, Table 3 summarizes claim-level diagnostics on the Filtered split. For LLaVA-2, 68% of claims are supported (hallucination 32%); ROUGE-L is 0.61. Faithfulness varies by domain, with WebQA higher ($\sim$72%) than Visual-RAG ($\sim$60%).
Hallucinations in Long Answers.
MM-RAGChecker shows that about one-third of answer statements are unsupported. Most are intrinsic hallucinations—plausible but unevidenced details (e.g., adding a material or breed not visible in the image)—while extrinsic hallucinations (off-topic filler) are rarer. For instance, in a Visual-RAG sample the model correctly states that a dog is catching a frisbee, but hallucinates its color and breed; our checker flags only those clauses.
Cross-modality behavior.
Beyond per-claim labels, we analyze modality interplay in Table 4, using $\DeltaCR/Δ\DeltaCP$ (image vs. text claim recall/precision gaps), VIS-Hit@$k$ and TXT-MissRate (evidence coverage), CMA (cross-modality agreement), and V-HR/D-HR (where hallucinations are mistakenly “entailment”). Large positive $\DeltaCR/Δ\DeltaCP$ indicates an image-leaning model, while high TXT-MissRate reveals wasted textual context. Lower CMA suggests the model tends to commit to a single modality instead of reconciling both.
Evidence Utilization and Modality Bias.
On average, the model uses $\sim$21% of retrieved pieces—suggesting poor integration. We observe a bias to stick to either text or image evidence instead of combining both, echoing prior findings that VLMs over-rely on text. Providing only gold evidence (GO mode) increases the Supported fraction to $\sim$90%, indicating distractors significantly hurt grounding.
Evidence Utilization.
Across datasets, LLaVA-2 uses on average 2.1 out of 10 retrieved items ($\sim$21% utilization), indicating under-use of available context. When provided only gold evidence (GO mode), the supported claim ratio climbs to about 90%, implying distractors directly harm grounding.
5.4 Effect of Retrieval Modes
Table 5 shows EM/Accuracy under DO, GPD, and GO. DO severely degrades performance (as expected), while GPD lies between DO and GO. The gap between GPD and GO quantifies the model’s susceptibility to distractors; Chart-RAG and VisRAG-Doc exhibit the largest drops, revealing brittle selection of evidence.
5.5 Prompt Sensitivity Study
We sweep 12 prompt designs (init style, few-shot exemplar count, reasoning pattern). Table 6 lists the best-performing prompt per dataset; Figure 3 visualizes the direct vs. retrieve-then-reason trend.
Table 6 reports, for each dataset, the prompt that achieves the highest mean Accuracy/EM across six models (filtered split). We observe clear patterns: Reasoning matters: retrieve_then_reason dominates on text-heavy retrieval tasks (MRAG, WebQA), whereas direct is consistently better on VisRAG-* visual QA splits where context is already focused. Few-shot count is task-dependent: Using three exemplars (ex3) suffices—and sometimes outperforms ex6—for purely visual/numeric QA (plots/slides). In contrast, ex6 helps more on document-style and open-web questions that benefit from pattern coverage. Init style is secondary: Switching between plain and expert seldom changes the rank unless combined with the right reasoning pattern and example count.
If a single prompt must be used across all datasets, expert-ex3-direct provides the best overall trade-off (top-1 on 3/6 datasets and close second elsewhere).
Dataset Best Prompt Mean Acc. (%)
MRAG-Bench expert3 retrieve & reason 31.5
VisRAG-ArXiv expert6 direct 5.0
VisRAG-Doc plain6 direct 34.3
VisRAG-Plot expert3 direct 36.5
VisRAG-Slide plain3 direct 35.2
WebQA expert6 retrieve&reason 23.0
Table 6: Best-performing prompt (by mean EM/Acc across models) per dataset on the filtered split. The patterns highlight when “retrieve-then-reason” helps (text-heavy tasks) versus when direct prompting is sufficient (focused visual tasks).
Figure 3:
5.6 Impact of Distractors on Diagnostic Metrics
[…]
5.7 Effect of Filtering
Filtering not only reduces accuracy but also changes answer style. Average long-answer length increases from 45 to 60 words after filtering, yet factuality does not improve—MM-RAGChecker flags many of the additional sentences as unsupported. This suggests that harder questions trigger verbosity rather than better grounding.
5.8 Takeaways
(1) Filtering and distractors expose large drops in both accuracy and grounding. (2) Current VLMs hallucinate frequently even when retrieval is successful; claim-level diagnostics are necessary. (3) Ambiguity is a major failure source: without explicit clarification, models either ignore alternative readings or fuse them incorrectly. (4) Improving evidence selection and utilization—possibly via tighter retrieval–generation coupling or in-the-loop verification—is a promising direction.
6 Conclusion
We presented MRAG-Suite, a comprehensive suite for evaluating and diagnosing visual retrieval-augmented generation in multimodal question answering. MRAG-Suite introduces a unified benchmark that stresses truly multimodal, knowledge-intensive QA, combining diverse sources like web images, textual articles, and informational charts. To focus evaluation on the intended challenges, we devised a two-step filtering method to remove trivial cases and emphasize multi-hop, vision+language reasoning. We also contributed an automated way to generate long-form answers with evidence, enabling novel evaluation of explanations in addition to factual answers. Crucially, we developed MM-RAGChecker, a claim-level multimodal verifier, to assess the factual consistency of generated answers.
Limitations
This work targets visual RAG for images, charts, slides, scanned pages, and scholarly PDFs; results may not transfer to videos, embodied settings, or low-resource languages. Hybrid retrieval depends on CLIP-like encoders and OCR text, so known dataset/model biases and OCR noise can affect measured recall/precision and inflate hallucination labels. Our diagnostics rely on LLM/VLM components for claim extraction and entailment; despite prompt controls and spot-checks, such judges can diverge from expert assessments in domain-specific cases. Scores are also sensitive to prompt format, retrieval $k$, and distractor sampling; we report settings but do not exhaust all combinations. Some baselines use third-party APIs, so provider updates and rate limits may introduce temporal variance despite fixed prompts and released scripts.
Appendix A Ambiguity Process
A.1 Ambiguity Definition and Annotation
We curate a 200-question subset to study ambiguity in multimodal RAG.
Definition.
A question is AMBIGUOUS if a reasonable reader cannot determine a single correct answer without additional constraints (time, location, unit, object, visibility in the image, etc.); otherwise it is CLEAR. For image-conditioned queries, we additionally mark as ambiguous when key visual evidence is not visible or is blurred.
Detection Pipeline.
We adopt a two-stage procedure:
  1. 1.
    LLM-based pre-filter. We prompt an assessor model to output CLEAR or AMBIGUOUS together with a short rationale (30 words). We use separate prompts for text-only and image-conditioned questions (see Appendix A.3).
  2. 2.
    Human adjudication. Two annotators independently review the flagged items and resolve disagreements. We achieve substantial agreement (Cohen’s $\kappa=0.74$).
Rewriting and Answer Normalization.
For each ambiguous item we:
  • Evidence-grounded answer rewrite: given retrieved evidence and the original answer, an LLM produces a concise answer or returns “Evidence inconclusive.” (Appendix A.5).
  • Query disambiguation: another prompt rewrites the question into a precise version that can be answered solely from the evidence (Appendix A.6).
The final ambiguity split (200 queries) and a matched clear set (200 queries) are used in Appendix. A.7 to isolate the impact of ambiguity on accuracy, claim recall, and hallucination rate.
A.2 Error Patterns under Ambiguity
On the 200-query ambiguity subset, two dominant patterns emerge:
  1. 1.
    Single-interpretation bias: the model commits to one reasonable sense, lowering Claim Recall and short-answer accuracy.
  2. 2.
    Interpretation conflation: retrieval returns evidence for multiple senses; the generator merges them, producing composite unsupported claims and higher hallucination rates.
These results motivate an ambiguity detection & clarification step before final answer generation.
A.3 Ambiguity Detection (Image-conditioned)
A.4 Ambiguity Detection (Text-only)
A.5 Evidence-based Answer Rewrite
A.6 Ambiguous Query Rewrite
Table 7: Ambiguity subset construction statistics. We report candidate counts, items flagged as ambiguous by the two-stage pipeline, and the final curated set (200).
Candidates Marked Ambiguous Final Set
Text-only 26,120 13,823 184
Image-conditioned 4,557 1,224 16
Total 30,677 15,047 200
A.7 Ambiguity Subset Results
On the 200-query ambiguous subset, short-answer EM drops by 5.9 points compared to clear queries, while hallucination rate increases by 9.5 points. Claim Recall and Context Precision also decrease, indicating single-interpretation bias or interpretation conflation; see Appendix. A.2 for qualitative patterns. Table 8 reports the aggregate metrics. Also, we have the ambiguous set in the code provided.
Table 8: Results on the 200-query ambiguity subset versus a matched clear set. We report EM/Accuracy, Hallucination Rate (lower is better), Claim Recall, and Context Precision, isolating the effect of query underspecification on retrieval and grounding.
EM/Acc $\uparrow$ Halluc. $\downarrow$ Claim Recall $\uparrow$ Ctx. Prec. $\uparrow$
Clear (200 random) 23.5 49.2 12.4 14.1
Ambiguous (200) 17.6 58.7 9.1 12.5
$\Delta$ (Ambig - Clear) -5.9 +9.5 -3.3 -1.6
Appendix B Retrieve system settings
Dataset Text Retriever Image Retriever
WebQA SBERT (mpnet) CLIP ViT-B/32
Chart-RAG DPR (ctx-enc) CLIP ViT-B/32
MRAG CLIP ViT-B/32 (img+txt avg)
VisRAG VisRAG-Ret
VisualRAG CLIP ViT-B/32
Table 9: Backbones used in the hybrid retriever for each sub-corpus. Images are indexed with CLIP variants; text uses DPR/SBERT; VisRAG documents use a document-image retriever.
Appendix C 12 Prompt Experiment Settings
Layer Arg name Values in this paper Template / Semantics
System Prompt init_style plain plain: “Please read the following question and retrieve the relevant document(s)/image(s) to produce the best possible answer.”
expert expert: “You are an expert in <DOMAIN>. Analyze the question carefully, think step-by-step about which sources to use, and present your answer clearly at the end.”
Few-shot Demo example_style ex1, ex3, ex6 Three fixed 0-shot style exemplars (chosen from the pool). Content differs in wording/format hints.
max_examples 0 or 5 Whether to insert 5 illustrative QA pairs (kept at 0 in our sweep).
Reasoning Instruction direct direct: “Question: … Answer:” (no mandated reasoning steps).
reasoning_process retrieve_then_reason: “First outline retrieval/reasoning, then give final answer.”
Answer Format answer_format free Free-form text (no forced JSON or letter).
Context Packaging context_order img_first Images precede text in the prompt (fixed in our runs).
include_doc_ids 0 / 1 Whether to show doc/image IDs (0 in our runs).
Table 10: Prompt dimensions and concrete templates used in our 12-way sweep.
Nickname init_style ex_style Example template
expert1 expert ex1 Q: {q}
A: {ans}
expert2 expert ex3 Question + brief Explanation + Final Answer
expert3 expert ex6 Dialog style: User: {q} / Assistant: {ans}
plain1 plain ex1 Q: {q}
A: {ans}
plain2 plain ex3 Question + brief Explanation + Final Answer
plain3 plain ex6 Dialog style: User: {q} / Assistant: {ans}
Table 11: Mapping from prompt nicknames to template families used in the 12-way prompt sweep. Variations differ in initialization style and exemplar formatting.
p_reason Reasoning block (excerpt) COT line?
direct Answer directly without explaining your reasoning. No
retrieve_then_reason First list which retrieved snippets/images you will use (by Doc/Image IDs). Then reason step by step using them, and finally give the answer. Yes
structured Follow this structure: 1) Question understanding 2) Relevant sources (IDs or short quotes) 3) Reasoning 4) Final Answer Yes
plan_execute Outline a short plan first (bullet points). After the plan, execute it and derive the final answer. Yes
verify Propose an initial answer. Then self-check it briefly. If it passes, output the final answer prefixed with “FINAL:”. Yes
none (no additional block) No
Table 12: Reasoning patterns injected into prompts. Variants differ in whether to expose chain-of-thought–like structure and explicit self-verification before the final answer.
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 Pixtral-12B-2409 InternVL3-8B
Filt Full Filt Full Filt Full Filt Full Filt Full Filt Full
Chart-MRAG 22 19 15 25 10 11 13 12 14 29 19 27
MRAG-Bench 32 33 19 24 28 26 23 24 38 37 21 21
VisRAG-ArXiv 55 48 34 34 35 35 26 27 43 43 63 55
VisRAG-Doc 31 35 20 21 28 29 24 22 26 26 39 34
VisRAG-Plot 36 45 12 13 35 36 6 9 51 56 33 36
VisRAG-Slide 18 25 8 10 17 22 24 22 18 15 20 22
Visual-RAG 13 13 9 11 11 10 8 9 12 13 11 12
WebQA 16 14 12 13 14 17 10 11 26 27 17 28
Table 13: EM/Accuracy performance (%) of each model on the Filtered and Full splits.
Appendix D Distractor results overview
Dataset Model Distractor Recall Precision F1 Halluc. Faith. Self-know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
VisRAG-ArXiv Qwen2.5-VL-72B distractors only 21.8 6.7 7.9 66.2 23.5 5.9 20.4 31.0
gt_plus_distractors 29.6 8.9 10.8 60.1 28.9 7.1 24.1 33.5
gt_only 33.4 10.8 12.7 55.3 32.1 8.2 25.9 35.8
LLaVa-v1.6-34B distractors only 18.4 5.9 6.8 69.4 20.1 4.8 18.0 28.2
gt_plus_distractors 24.7 7.6 9.1 63.5 24.9 5.9 21.6 30.4
gt_only 28.9 9.2 10.9 58.2 29.6 6.8 23.7 33.1
Claude 3.5 distractors only 30.4 9.2 10.1 61.4 28.0 7.6 29.9 40.0
gt_plus_distractors 40.7 10.6 14.3 56.0 37.0 7.0 25.6 33.0
gt_only 35.9 8.6 11.3 57.7 36.5 4.7 27.6 39.0
Phi-4 distractors only 22.5 15.3 13.7 50.8 20.6 12.9 22.2 33.8
gt_plus_distractors 34.4 24.0 23.5 45.3 22.3 16.3 26.0 36.0
gt_only 37.8 27.2 26.6 41.1 24.7 18.5 27.9 37.9
Pixtral-12B-2409 distractors only 17.6 7.1 6.3 70.5 18.7 5.1 16.0 27.6
gt_plus_distractors 25.9 10.4 9.7 63.0 23.3 7.4 18.9 29.5
gt_only 30.5 12.8 12.0 58.9 25.6 8.3 21.5 31.2
InternVL3-8B distractors only 26.5 14.6 12.6 59.3 24.0 10.8 28.8 37.0
gt_plus_distractors 39.2 15.7 16.5 54.4 30.0 10.6 29.0 38.0
gt_only 40.3 14.6 17.6 48.2 27.4 9.5 24.4 35.0
VisRAG-Doc Qwen2.5-VL-72B distractors only 10.8 4.0 3.6 57.9 13.4 2.6 21.0 23.5
gt_plus_distractors 46.7 31.9 27.2 28.1 22.7 20.4 25.9 27.0
gt_only 52.9 36.2 31.0 24.4 24.5 22.6 24.1 25.1
LLaVa-v1.6-34B distractors only 9.4 3.7 3.2 60.1 12.1 2.3 19.5 22.0
gt_plus_distractors 39.8 26.1 22.0 33.9 20.1 16.4 23.8 25.6
gt_only 44.0 28.9 24.7 30.2 22.4 18.5 22.9 24.2
Claude 3.5 distractors only 12.0 4.6 4.2 38.2 16.9 3.0 24.5 26.0
gt_plus_distractors 75.5 35.2 38.2 41.4 32.2 25.4 27.7 31.0
gt_only 68.8 33.6 33.7 41.5 30.7 23.8 26.5 28.0
Phi-4 distractors only 18.1 12.2 11.0 27.6 7.5 10.5 20.8 22.7
gt_plus_distractors 45.3 37.8 37.0 15.2 4.7 32.3 24.5 26.0
gt_only 50.6 41.9 41.1 13.0 3.9 35.2 23.6 24.8
Pixtral-12B-2409 distractors only 7.8 3.5 3.0 61.3 10.6 2.1 18.8 21.3
gt_plus_distractors 34.4 25.6 20.4 36.8 17.7 15.1 22.1 23.9
gt_only 39.1 29.7 24.1 32.6 19.8 17.2 21.5 22.8
InternVL3-8B distractors only 8.0 4.1 3.4 55.2 12.8 3.1 22.5 25.0
gt_plus_distractors 68.3 65.2 26.5 15.8 25.6 45.5 28.0 26.5
gt_only 60.3 77.8 57.8 17.2 27.3 54.5 21.5 23.0
MRAG-Bench Qwen2.5-VL-72B distractors only 22.6 6.2 5.1 79.4 14.3 4.0 7.3 4.5
gt_plus_distractors 28.4 7.9 6.7 75.8 16.2 4.7 8.9 5.2
gt_only 33.7 9.5 8.4 71.1 18.6 5.3 10.4 6.0
LLaVa-v1.6-34B distractors only 21.1 5.3 4.1 82.4 13.2 3.5 8.2 4.2
gt_plus_distractors 27.8 4.8 3.1 77.1 17.6 4.3 10.1 6.0
gt_only 31.9 6.7 4.6 73.4 19.1 4.8 9.8 5.0
Claude 3.5 distractors only 42.2 3.9 4.6 73.5 19.8 3.8 9.2 5.5
gt_plus_distractors 47.2 2.8 3.9 80.5 16.1 2.4 7.7 6.0
gt_only 47.7 4.2 5.6 78.8 17.1 4.1 5.5 3.0
Phi-4 distractors only 14.5 12.6 10.8 36.4 8.8 11.9 7.1 4.6
gt_plus_distractors 18.7 18.3 16.2 31.4 9.3 16.3 8.6 5.5
gt_only 22.6 21.4 18.7 27.8 10.4 18.5 9.5 6.2
Pixtral-12B-2409 distractors only 12.8 6.9 5.7 84.6 11.0 5.6 6.4 3.9
gt_plus_distractors 17.2 8.8 7.2 80.3 12.5 6.4 7.3 4.2
gt_only 21.0 10.6 8.6 76.1 14.3 7.2 8.4 4.9
InternVL3-8B distractors only 35.3 37.6 31.0 54.5 9.3 35.1 5.3 3.0
gt_plus_distractors 29.0 33.1 25.1 46.7 8.2 32.1 5.3 3.5
gt_only 31.8 31.6 25.6 62.0 7.5 28.6 6.2 4.0
VisRAG-Plot Qwen2.5-VL-72B distractors only 10.1 3.9 2.6 61.7 12.2 2.7 9.4 10.2
gt_plus_distractors 17.6 7.2 5.8 58.5 11.6 5.4 9.7 9.6
gt_only 23.8 10.7 8.2 56.0 12.9 7.7 8.6 8.8
LLaVa-v1.6-34B distractors only 7.9 3.1 1.9 64.9 10.7 2.1 9.1 10.4
gt_plus_distractors 13.4 6.1 4.0 61.2 11.5 3.9 9.5 9.8
gt_only 19.6 8.3 5.7 58.9 12.1 5.5 8.9 9.3
Claude 3.5 distractors only 8.0 4.1 2.1 55.1 10.1 2.8 10.5 11.0
gt_plus_distractors 19.8 9.4 6.6 77.8 13.4 8.9 9.0 9.0
gt_only 26.5 11.6 9.4 75.2 12.7 11.1 7.5 8.0
Phi-4 distractors only 11.3 4.7 1.8 83.7 10.3 4.0 10.0 10.0
gt_plus_distractors 16.8 8.9 3.7 79.5 11.2 6.8 9.6 9.7
gt_only 22.4 11.5 5.2 76.4 12.0 9.4 8.4 8.7
Pixtral-12B-2409 distractors only 6.5 2.8 1.1 66.0 9.8 1.7 8.8 9.5
gt_plus_distractors 12.9 6.4 2.9 62.1 10.5 3.7 9.2 9.6
gt_only 18.7 8.9 4.6 59.6 11.2 5.2 8.7 9.0
InternVL3-8B distractors only 2.3 1.8 0.4 52.3 11.0 1.7 8.5 9.0
gt_plus_distractors 25.5 15.1 7.9 72.9 9.3 13.8 10.0 10.0
gt_only 21.5 11.2 3.2 76.5 8.2 10.3 11.1 12.0
VisRAG-Slide Qwen2.5-VL-72B distractors only 12.6 5.5 4.8 63.3 14.7 3.8 51.1 56.8
gt_plus_distractors 41.3 27.2 26.1 36.5 42.2 21.1 46.9 50.6
gt_only 49.7 34.9 33.5 31.2 48.4 23.2 43.8 48.0
LLaVa-v1.6-34B distractors only 10.9 4.9 4.1 66.0 13.3 3.3 49.6 54.3
gt_plus_distractors 35.6 22.7 21.0 40.2 39.1 18.8 45.1 49.2
gt_only 42.8 29.6 27.4 35.5 43.7 20.6 41.9 46.6
Claude 3.5 distractors only 23.1 12.0 11.2 53.3 36.7 9.1 53.3 59.1
gt_plus_distractors 78.8 30.6 37.3 38.7 45.8 15.4 55.7 61.1
gt_only 64.2 39.1 36.0 30.1 49.2 18.9 49.2 56.0
Phi-4 distractors only 19.9 11.7 10.3 41.2 31.6 8.6 50.1 55.6
gt_plus_distractors 71.8 42.7 44.5 33.7 40.1 23.4 45.9 48.6
gt_only 60.4 47.9 45.0 29.0 37.7 26.1 43.6 46.2
Pixtral-12B-2409 distractors only 9.6 5.1 4.2 67.9 12.0 3.4 48.5 53.7
gt_plus_distractors 29.8 18.9 16.7 45.6 34.7 16.3 44.3 47.5
gt_only 36.7 24.5 22.1 39.8 38.9 17.9 41.1 45.3
InternVL3-8B distractors only 17.2 4.4 5.1 58.7 34.0 3.2 54.1 59.7
gt_plus_distractors 67.1 42.6 41.9 23.4 51.4 22.5 47.3 51.4
gt_only 57.8 41.1 38.7 28.8 44.6 23.8 40.8 45.8
WebQA Qwen2.5-VL-72B distractors only 14.0/13.1 6.5/5.8 5.0/4.3 69.1/71.2 9.4/8.6 5.0/5.4 12.8/12.0 14.0/13.2
gt_plus_distractors 18.3/17.2 8.9/8.1 7.3/6.5 66.5/68.7 11.2/10.3 6.3/6.8 12.0/11.3 13.5/12.6
gt_only 22.4/21.0 11.3/10.2 9.1/8.0 62.8/64.9 13.0/12.0 7.4/8.0 11.2/10.5 12.7/11.8
LLaVa-v1.6-34B distractors only 12.6/11.4 5.7/5.0 4.5/3.8 71.6/73.9 8.0/7.2 4.8/5.1 12.1/11.3 13.1/12.1
gt_plus_distractors 16.8/15.6 8.2/7.4 6.6/5.7 68.7/70.8 10.4/9.5 5.9/6.3 11.3/10.6 12.7/11.8
gt_only 21.3/19.9 10.5/9.4 8.4/7.3 65.2/67.3 12.6/11.6 7.0/7.6 10.7/10.0 12.0/11.2
Claude 3.5 distractors only 23.0/21.5 4.7/4.3 5.2/4.4 63.4/86.4 31.9/4.1 3.7/3.5 12.6/4.8 15.0/12.0
gt_plus_distractors 22.4/20.7 5.9/4.3 5.4/4.7 61.1/87.2 30.9/4.4 4.9/3.3 10.1/7.0 13.0/16.5
gt_only 26.8/24.1 6.8/5.0 6.6/5.4 57.6/84.3 33.5/5.2 5.5/3.9 11.2/8.1 14.2/17.8
Phi-4 distractors only 14.2/12.7 22.0/19.1 8.7/7.5 53.4/60.2 9.1/0.7 20.1/18.5 10.9/4.7 13.6/12.3
gt_plus_distractors 10.4/12.5 25.6/22.3 9.5/11.1 49.7/56.5 8.4/0.5 23.9/22.0 11.8/5.2 14.5/13.0
gt_only 16.1/14.8 29.0/24.7 12.1/12.4 46.2/53.0 10.2/0.9 26.7/24.1 12.3/5.8 15.9/13.9
Pixtral-12B-2409 distractors only 10.5/ 9.6 15.3/14.1 6.4/5.7 58.9/65.2 6.9/1.0 13.8/12.6 9.6/3.9 12.0/10.9
gt_plus_distractors 13.4/12.6 19.6/17.4 8.7/7.5 55.2/61.8 8.0/1.2 17.3/15.5 10.5/4.4 13.2/11.9
gt_only 17.4/16.0 23.5/20.7 11.3/9.8 51.7/58.1 9.6/1.6 20.4/18.0 11.8/5.0 14.7/13.0
InternVL3-8B distractors only 27.9/27.3 13.8/14.6 14.7/12.8 68.1/77.2 13.8/8.8 13.1/11.9 13.0/4.5 14.0/10.5
gt_plus_distractors 32.0/31.0 15.4/15.3 17.0/15.1 69.0/74.4 14.2/8.4 14.8/12.2 11.2/6.3 14.0/14.0
gt_only 16.3/16.9 7.0/6.9 5.6/5.1 64.0/63.6 8.9/12.0 6.1/5.3 8.5/6.4 11.5/13.5
Table 14: Full $M$-RAGChecker diagnostics across datasets, models, and retrieval modes. For each setting we report per-modality (image/text) Recall, Precision, F1, Hallucination, Faithfulness, Self-Knowledge, Claim Recall, and Context Precision.
Dataset Model Distractor Recall Precision F1 Halluc. Faith. Self‑know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
ChartMRAG-Bench Qwen2.5-VL-72B distractors only 11.0/10.0 4.2/3.7 2.9/2.4 60.6/62.8 12.8/11.9 2.6/2.9 9.8/9.1 10.6/10.0
gt_plus_distractors 18.4/17.1 7.6/6.9 6.2/5.4 57.1/59.4 12.2/11.1 5.1/5.6 10.1/9.4 10.0/9.7
gt_only 24.6/23.1 11.1/10.3 8.6/7.7 54.6/56.9 13.4/12.6 7.4/8.1 9.0/8.3 9.2/8.9
LLaVa-v1.6-34B distractors only 8.6/7.6 3.4/2.9 2.2/1.7 63.7/65.8 11.2/10.5 2.0/2.3 9.5/8.8 10.7/10.2
gt_plus_distractors 14.2/13.0 6.5/5.8 4.3/3.5 60.1/62.3 12.0/11.0 3.8/4.2 9.9/9.2 10.2/9.8
gt_only 20.7/19.2 8.9/8.1 6.1/5.3 57.8/60.1 12.6/11.8 5.4/6.0 9.2/8.6 9.6/9.1
Claude 3.5 distractors only 8.6/7.5 4.3/3.9 2.3/1.9 54.0/56.0 10.4/9.8 2.7/3.0 10.8/10.2 11.2/10.7
gt_plus_distractors 20.7/18.9 9.8/8.9 6.9/6.2 76.6/78.9 13.8/12.7 9.2/9.8 9.2/8.9 9.2/8.8
gt_only 27.5/25.4 12.2/11.0 9.7/8.9 74.3/76.6 13.1/12.2 11.2/11.9 7.7/7.3 8.2/7.8
Phi-4 distractors only 12.1/10.6 4.9/4.4 2.0/1.7 82.4/84.7 10.9/9.8 3.7/4.0 10.2/9.7 10.2/9.9
gt_plus_distractors 17.6/16.0 9.4/8.5 4.1/3.4 78.3/80.6 11.7/10.8 6.5/7.0 9.9/9.3 9.9/9.5
gt_only 23.3/21.6 12.0/10.9 5.5/4.9 75.1/77.4 12.5/11.6 9.1/9.7 8.7/8.1 8.9/8.5
Pixtral-12B-2409 distractors only 7.2/6.0 3.1/2.6 1.3/1.0 64.8/66.9 10.3/9.4 1.6/1.9 9.1/8.5 9.8/9.3
gt_plus_distractors 13.7/12.2 6.7/5.9 3.1/2.7 60.9/63.1 11.0/10.1 3.6/3.9 9.5/8.9 9.9/9.4
gt_only 19.6/18.0 9.3/8.2 4.9/4.3 58.6/60.8 11.7/10.8 5.1/5.6 9.1/8.5 9.3/8.9
InternVL3-8B distractors only 2.6/2.1 1.9/1.6 0.5/0.4 51.1/53.5 11.3/10.7 1.6/1.8 8.7/8.3 9.2/8.8
gt_plus_distractors 26.5/24.6 15.7/14.4 8.4/7.5 71.8/74.0 9.7/8.9 14.3/15.0 10.1/9.7 10.1/9.8
gt_only 22.3/20.8 11.6/10.7 3.5/2.9 75.4/77.5 8.5/7.8 10.6/11.1 11.4/10.8 12.1/11.4
Visual-RAG Qwen2.5‑VL‑72B distractors only 18.4 6.3 6.8 42.1 27.2 48.3 4.9 18.2
gt_plus_distractors 44.5 14.7 16.8 36.3 34.5 53.9 6.8 25.0
gt_only 41.3 13.2 15.4 37.5 32.8 51.1 6.5 23.3
LLaVa‑v1.6‑34B distractors only 16.1 5.2 5.7 43.7 25.4 45.0 4.4 16.1
gt_plus_distractors 42.0 12.9 14.7 38.8 30.1 51.6 6.0 22.7
gt_only 38.8 11.8 13.2 39.5 28.0 50.2 5.7 21.4
Claude 3.5 distractors only 23.3 7.0 8.4 40.2 28.6 50.3 6.2 20.0
gt_plus_distractors 52.7 16.5 19.7 35.0 36.7 55.5 7.5 28.8
gt_only 48.1 15.2 17.4 36.1 34.9 53.2 7.1 26.5
Phi‑4 distractors only 20.7 6.5 7.3 39.0 26.0 46.4 5.3 18.0
gt_plus_distractors 47.2 15.0 17.6 34.1 32.5 52.9 6.7 25.2
gt_only 43.6 13.9 16.0 35.3 31.0 50.7 6.4 23.7
Pixtral‑12B‑2409 distractors only 15.3 5.8 6.2 44.9 23.5 43.2 4.1 15.3
gt_plus_distractors 40.8 12.0 13.9 38.2 29.8 49.8 5.8 21.9
gt_only 37.9 11.1 12.8 39.0 27.7 48.1 5.5 20.4
InternVL3‑8B distractors only 22.6 8.1 9.2 41.3 30.0 49.0 6.0 19.5
gt_plus_distractors 50.1 17.2 20.3 33.7 37.8 54.4 7.8 27.1
gt_only 45.7 15.5 18.1 34.9 35.3 52.0 7.3 25.0
Table 15: