### v2

# MRAG-Suite: A Cross-Domain Diagnostic Benchmark and Tooling for Visual Retrieval-Augmented Generation

Abstract
We introduce MRAG-Suite, a cross-domain evaluation suite for Visual Retrieval-Augmented Generation (RAG) that stresses real-world retrieval, multimodal grounding, and robustness. The suite consolidates eight heterogeneous sources (web photos, charts/plots, scanned documents, slides, scholarly PDFs) under unified formats and metrics. To better reflect realistic usage, we add two-step difficulty filtering and a 200-item ambiguity split, and evaluate with three retrieval modes (gold-only, gold+ distractors, distractors-only). We also present MM-RAGChecker, a claim-level diagnostic judge for multimodal grounding, reporting hallucination, faithfulness, claim recall, context precision, and cross-modality agreement/attribution. Across open-source and proprietary VLMs, we observe large drops under difficult/ambiguous queries and in the presence of distractors, as well as strong modality biases and under-utilization of retrieved evidence. MRAG-Suite and MM-RAGChecker provide actionable, fine-grained signals for improving evidence selection and justification in visual RAG systems.
MRAG-Suite: A Cross-Domain Diagnostic Benchmark and Tooling for Visual Retrieval-Augmented Generation
Yuelyu Ji University of Pittsburgh yueluji@gmail.com          Patrick NG patrickldld@gmail.com
1 Introduction
Retrieval-augmented generation (RAG) grounds large language models (LLMs) in external evidence to improve factuality and reliability. With the rise of multimodal models, visual RAG—jointly leveraging text and images—has gained traction for complex questions involving charts, plots, figures, photos, and document pages. Yet evaluation has not kept pace: most multimodal benchmarks either assume pre-selected visuals (thus bypassing retrieval) or remain text-only, limiting our ability to measure retrieval accuracy, multimodal grounding, and robustness in realistic settings.
Existing multimodal QA datasets such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and MMMU-Pro (Yue2024MMMUProAM) supply the necessary visual context in the input, side-stepping retrieval. Retrieval-centric efforts like Visual Haystacks (Wu2024VisualHA) emphasize hard look-up but focus on a single domain and lack fine-grained diagnosis. Text-focused diagnostic tools (e.g., RAGChecker, RefChecker) (ru2024ragchecker; Hu2024RefCheckerRF) provide claim-level analysis, but cannot assess how claims are grounded in visual evidence.
Figure 1: Illustration of our multimodal claim verification process. Given a long-form answer from a RAG system, we extract atomic claims and assess support from retrieved visual evidence. Claims are labeled Entailment if at least one image supports them. In this example, several traits are correctly described, but the gold claim about “high-altitude environments” remains unsupported.
To bridge these gaps, we introduce MRAG-Suite, a cross-domain diagnostic evaluation suite for visual RAG. MRAG-Suite integrates eight heterogeneous sources spanning charts/plots, natural photos, scanned documents, slides, and scholarly PDFs, and standardizes them under unified formats and metrics. To reflect real-world usage, the suite (i) applies difficulty-based filtering to de-trivialize questions, (ii) builds a 200-item ambiguity split to test underspecified queries, and (iii) evaluates three retrieval modes (gold-only, gold+ distractors, distractors-only) to probe robustness against distractors.
We further present MM-RAGChecker, a claim-level multimodal judge that verifies long-form answers against retrieved text and images. It reports hallucination and faithfulness alongside claim recall, context precision, and cross-modality behavior (e.g., agreement and attribution), yielding actionable diagnostics for error analysis and model development.
Contributions.
  1. 1.
    A unified, multi-domain visual RAG benchmark. We consolidate and normalize datasets across diverse domains, providing unified evaluation standards and enabling direct comparisons of model performance across various multimodal tasks.
  2. 2.
    Difficulty- and ambiguity-aware evaluation. Two-step filtering removes trivial items; an ambiguity subset stresses underspecified queries. Controlled distractor settings quantify robustness.
  3. 3.
    Claim-level multimodal diagnostics. MM-RAGChecker extends textual claim checking (ru2024ragchecker; Hu2024RefCheckerRF) to visual contexts, attributing support or errors to specific modalities and retrieved items.
All code, prompts, and diagnostics are available at:
https://anonymous.4open.science/status/MRAGChecker-B33D
2 Related Work
Retrieval-Augmented Generation Evaluation
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (ru2024ragchecker) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
Multimodal QA and Reasoning Benchmarks
Benchmarks such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and their advanced variants MMMU-Pro (Yue2024MMMUProAM) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu2024VisualHA) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsli2025human; li2025chatmotion.
Biases in Multimodal Retrieval-Augmented Systems
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao2025WhoII) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlywang2025selfdestructivelanguagemodel; liang2025graphrag; jiang2024robustkv; dan2024evaluation; lu2025predicting; li2025frequency.
Domain-specific Multimodal QA
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (yang2025benchmarking) and VDocRAG (Tanaka2025VDocRAGRG), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
3 MRAG-Suite Benchmark Construction
3.1 Task Definition
We evaluate multimodal Retrieval-Augmented Generation (RAG) for open-ended QA. Given a query $q$, the system retrieves textual ($E^{\text{txt}}$) and visual ($E^{\text{img}}$) evidence and must produce (i) a concise short answer and (ii) a grounded long answer: $(a^{\text{short}},a^{\text{long}})=f\!\big(q,E^{\text{txt}},E^{\text{img}}\big).$
3.2 Corpora Overview
MRAG-Suite merges eight heterogeneous sources spanning charts, natural photos, open-web images, scanned PDFs, and slides. Each source contributes distinct modalities and reasoning demands:
  • WebQA chang2022WebQA: open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG yang2025benchmarking(from ChartQA masry2022chartqa): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu2025VisualRAGBT: queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench hu2024mrag: 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu2024VisRAGVR subsets: Arxiv Li2024MultimodalAA (scholarly PDF pages), Plot Methani2019PlotQARO (technical plots), Slide 10.1609/aaai.v37i11.26598 (slide decks), and Document Tito2022HierarchicalMT (scanned manuals/forms). Each page is treated as a visual document with OCR text.
Dataset statistics.
Table 1 summarizes domains, filtering counts, image resolutions, and text length statistics.
Table 1: Overview of the eight sources in MRAG-Suite, covering domains, filtering counts, image resolution, and question/answer lengths. The suite spans open-web images, charts/plots, scanned documents, slides, and scholarly PDFs to enable cross-domain visual RAG evaluation.
Dataset Domain Two-way filter Avg. Img (px) Max Q Max A Avg Q Avg A Ambig.#
[…]
Figure 2: $M$-RAGChecker overview: claim extraction, multimodal evidence matching, and verification with per-claim labels and diagnostic metrics.
3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captionsradford2021learning; text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.
3.4 Two-Step Filtering for Non-triviality
We remove trivial and guessable items via:
  1. 1.
    Retrieval-Independent Filter: questions solved by a strong closed-book model (Claude 3.5, confidence >0.9>0.9) or whose answers appear verbatim in the question.
  2. 2.
    Difficulty-Based Filter: rank remaining items by multi-hop requirement, modality dependency, and baseline success rate; drop the easiest $\sim$10% per domain.
We release two query variants: Filtered (minimal, disambiguated) and Full (original, potentially noisy).
Ambiguity-aware evaluation.
To isolate the impact of query ambiguity on Visual RAG, we construct an ambiguity-aware subset (200 items) using a two-stage pipeline: an LLM-based pre-filter with short rationales, followed by double-annotator adjudication (substantial agreement, $\kappa\!=\!0.74$). We also provide evidence-grounded answer rewriting and disambiguated query rewrites for these items. Full prompts and guidelines are deferred to Appendix A.1.
3.5 Result of Distractors
To probe robustness, we evaluate three settings:
  • gt_only (GO): only gold evidence;
  • gt_plus_distractors (GPD): gold mixed with CLIP-similar distractors;
  • distractors_only (DO): only distractors, testing spurious reasoning.
3.6 Output Specification
Systems must return both short and long answers. We compute EM/Accuracy for $a^{\text{short}}$, and claim-level diagnostics plus ROUGE-L for $a^{\text{long}}$ (MM-RAGChecker in Sec. 4).
Having standardized data and retrieval settings, we next introduce MM-RAGChecker, a claim-level diagnostic tool that evaluates how long-form answers are grounded in the retrieved multimodal evidence. It operates in three stages—claim extraction, evidence matching, and verification—and produces a set of fine-grained metrics for hallucination, faithfulness, and cross-modality behavior.
Evaluation axes (metric families).
We evaluate along two complementary metric families rather than by answer length: (i) End-task accuracy on the short answer ($a^{\text{short}}$) via EM/Acc; and (ii) Claim-level diagnostics on the long answer ($a^{\text{long}}$) using MM-RAGChecker (§4; Hallucination, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, and cross-modality metrics).
4 MM-RAGChecker: Multimodal Claim-Level Diagnosis
4.1 Overview
Stage 1 — Claim extraction. We split the long answer $a^{\text{long}}$ into minimal verifiable units $\mathcal{C}=\{c_{i}\}$ using simple heuristics plus an LLM splitter for compound clauses (e.g., numeric comparisons).
Stage 2 — Per-evidence judging (3-way). Given images $\{I_{k}\}_{k=1}^{K_{\mathrm{img}}}$ and text passages $\{t_{k}\}_{k=1}^{K_{\mathrm{txt}}}$ retrieved for the question, we directly judge each pair $(c_{i},e)$, $e\in\{I_{k}\}\cup\{t_{k}\}$, with a three-way decision $\{\textsc{Entailment},\textsc{Neutral},\textsc{Contradiction}\}$. No image captions are synthesized and no numeric scores are used. We cap to top $K_{\mathrm{img}}{=}3$ and $K_{\mathrm{txt}}{=}3$ by default and set temperature to zero for determinism. For each claim we store two multisets of judgments $\mathcal{J}^{\mathrm{img}}_{i}=\{j(c_{i},I_{k})\}$ and $\mathcal{J}^{\mathrm{txt}}_{i}=\{j(c_{i},t_{k})\}$.
Stage 3 — Aggregation to claim labels. Each claim receives a single final label $L_{i}\in\{\textsc{Entailment},\textsc{Neutral},\textsc{Contradiction}\}$ via a simple precedence rule: (i) $L_{i}{=}\textsc{Entailment}$ if any $j\in\mathcal{J}^{\mathrm{img}}_{i}\cup\mathcal{J}^{\mathrm{txt}}_{i}$ is Entailment; (ii) else $L_{i}{=}\textsc{Contradiction}$ if any judgment is Contradiction; (iii) else $L_{i}{=}\textsc{Neutral}$. For modality analysis, define binary flags $s^{\mathrm{img}}_{i}{=}1$ iff some $j\in\mathcal{J}^{\mathrm{img}}_{i}$ is Entailment, and analogously $s^{\mathrm{txt}}_{i}$ for text. We also mark a retrieved item $e$ as used if there exists a claim $c_{i}$ with $j(c_{i},e){=}\textsc{Entailment}$. All metrics below are computed on these final labels and flags.
4.2 Metrics
Core metrics.
Let $\mathcal{C}$ be the set of extracted claims, $L_{i}$ the final label of $c_{i}$, and $\mathcal{E}$ the set of retrieved items. Then
$\mathrm{Hallucination\ Rate}$ $=\frac{\#\{\,i:L_{i}=\textsc{Neutral}\,\}}{|\mathcal{C}|},$ (1)
$\mathrm{Faithfulness}$ $=\frac{\#\{\,i:L_{i}=\textsc{Entailment}\,\}}{|\mathcal{C}|}.$ (2)
Let $\mathcal{G}$ be gold salient claims mined from a reference long answer.
$\mathrm{Claim\ Recall\ (CR)}$ $=\frac{\#\{\text{gold }g\in\mathcal{G}\ \text{that appear in }a^{\text{long}}\}}{|\mathcal{G}|}.$ (3)
Let $\mathcal{E}_{\mathrm{used}}=\{e\in\mathcal{E}:\exists i,\ j(c_{i},e)=\textsc{Entailment}\}$.
$\mathrm{Context\ Precision\ (CP)}$ $=\tfrac{|\mathcal{E}_{\mathrm{used}}|}{|\mathcal{E}|}.$ (4)
Self-knowledge measures how many entailed claims did not rely on retrieval:
$\mathrm{Self\text{-}Knowledge}=\frac{\#\{\textsc{Entailment }c_{i}\ \text{with no }j(c_{i},e)=\textsc{Entailment}\ \forall e\in\mathcal{E}\}}{\#\{\textsc{Entailment }c_{i}\}}$
Cross-modality metrics.
We quantify cross-modality behavior with: Claim Recall (CR), Claim Precision (CP), Visual Hit@$k$ (VIS-Hit@$k$), Text Miss Rate (TXT-MissRate), Cross-Modal Agreement (CMA), Vision-attributed Hallucination Rate (V-HR), and Document/Text-attributed Hallucination Rate (D-HR). Let $\mathcal{C}$ be the set of generated atomic claims for a sample and $L_{i}\in\{\textsc{Entailment},\textsc{Contradicted},\textsc{Neutral}\}$ the verdict from MM-RAGChecker for claim $i\in\mathcal{C}$ with respect to the provided context. Let $\mathcal{I}$ and $\mathcal{T}$ denote the retrieved image and text sets; we write $v\models i$ if evidence $v$ entails claim $i$.
$s^{\mathrm{img}}_{i}$ $=\mathbf{1}\{\,\exists v\in\mathcal{I}:v\models i\,\},$ (6)
$s^{\mathrm{txt}}_{i}$ $=\mathbf{1}\{\,\exists t\in\mathcal{T}:t\models i\,\}.$ (7)
Claim-level recall/precision for $m\in\{\mathrm{img},\mathrm{txt}\}$:
[…]
$\mathrm{CP}_{m}$ $=\frac{\sum_{i\in\mathcal{C}}\mathbf{1}[s^{m}_{i}=1]}{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1\,]}.$ (9)
Cross-modality gaps:
$\Delta\mathrm{CR}$ $=\mathrm{CR}_{\mathrm{img}}-\mathrm{CR}_{\mathrm{txt}},$ (10)
$\Delta\mathrm{CP}$ $=\mathrm{CP}_{\mathrm{img}}-\mathrm{CP}_{\mathrm{txt}}.$ (11)
Coverage (restricting to top-$k$ per modality):
$\mathrm{VIS\!-\!Hit@}k$ $=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}\!\big[\,I_{i}\vDash\exists c\,\big],$ (12)
$\mathrm{TXT\!-\!MissRate}$ $=1-\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}\!\big[\,t_{i}\vDash\exists c\,\big].$ (13)
Cross-modal agreement:
$\mathrm{CMA}=\frac{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\wedge s^{\mathrm{txt}}_{i}=1\,]}{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1\,]}.$ (14)
Hallucination attribution:
$\mathrm{V\!-\!HR}$ $=\frac{\#\{\,i:\ s^{\mathrm{img}}_{i}=1\ \wedge\ L_{i}\neq\textsc{Entailment}\,\}}{\#\{\,i:\ L_{i}\neq\textsc{Entailment}\,\}},$ (15)
$\mathrm{D\!-\!HR}$ $=\frac{\#\{\,i:\ s^{\mathrm{txt}}_{i}=1\ \wedge\ L_{i}\neq\textsc{Entailment}\,\}}{\#\{\,i:\ L_{i}\neq\textsc{Entailment}\,\}}.$ (16)
Unless noted, all values are percentage points on the Filtered split, and we use $k=5$ per modality at test time.
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 Pixtral-12B-2409 InternVL3-8B
Filt Full Filt Full Filt Full Filt Full Filt Full Filt Full
Chart-MRAG 22.4 19.1 15.3 25.7 10.8 11.9 11.2 13.5 14.7 29.4 19.2 27.6
MRAG-Bench 32.6 33.8 19.5 24.6 28.3 26.1 32.4 30.7 38.8 37.5 21.9 21.4
VisRAG-ArXiv 55.3 48.7 34.6 34.5 35.5 35.1 38.6 36.9 43.1 43.7 63.9 55.4
VisRAG-Doc 31.1 35.4 20.2 21.9 28.5 29.2 27.6 28.2 26.9 26.6 39.7 34.6
VisRAG-Plot 36.7 45.8 12.3 13.6 35.9 36.8 6.4 9.2 51.4 56.5 33.5 36.1
VisRAG-Slide 18.5 25.3 8.7 10.4 17.6 22.1 24.3 22.6 18.2 15.5 20.8 22.9
Visual-RAG 13.8 13.1 9.6 11.5 11.7 10.6 8.5 9.2 12.1 13.3 11.2 12.6
WebQA 16.2 14.9 12.7 13.4 14.8 17.3 20.6 22.1 26.4 27.9 17.8 28.7
Table 2: EM/Accuracy performance of each model on the Filtered (retained valid samples) and Full (all samples) splits. Best and Second-best are highlighted per row and split.
Dataset Model Recall Precision F1 Halluc. Faith. Self-know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
Chart-MRAG Qwen2.5-VL-72B 12.1/12.9 7.9/8.1 6.3/7.2 57.3/58.9 54.9/7.8 8.3/7.7 8.9/14.7 85.2/31.8
LLaVa-v1.6-34B 7.6/10.1 0.6/18.9 1.1/13.2 57.4/66.8 50.5/8.8 3.6/10.6 2.0/12.8 86.1/22.3
Claude 3.5 13.3/13.2 0.8/2.1 1.4/3.4 24.9/87.3 65.1/10.6 0.0/2.1 75.0/10.0 90.0/15.0
MRAG-Bench Qwen2.5-VL-72B 16.3 1.2 2.2 45.2 62.6 12.4 11.3 72.1
LLaVa-v1.6-34B 10.9 1.0 1.8 57.8 54.8 3.6 8.5 87.4
Claude 3.5 53.3 4.1 7.6 48.2 66.3 5.7 11.0 79.6
VisRAG-ArXiv Qwen2.5-VL-72B 23.7 1.9 3.5 30.3 67.2 17.2 23.2 61.0
LLaVa-v1.6-34B 19.7 1.7 3.1 37.0 65.5 12.3 19.4 67.4
Claude 3.5 17.0 1.6 2.9 41.4 64.4 9.1 16.9 71.6
VisRAG-Doc Qwen2.5-VL-72B 16.8 1.5 2.8 45.0 57.5 8.0 11.1 74.3
LLaVa-v1.6-34B 11.0 0.9 1.6 54.3 56.3 4.2 5.6 82.3
Claude 3.5 12.1 1.2 2.2 51.6 57.7 4.7 7.3 80.4
VisRAG-Plot Qwen2.5-VL-72B 17.7 1.9 3.4 42.1 65.5 9.6 16.4 75.1
LLaVa-v1.6-34B 9.5 0.8 1.5 60.9 50.1 3.2 1.9 88.6
Claude 3.5 17.1 1.5 2.8 46.7 62.8 8.3 13.5 75.9
VisRAG-Slide Qwen2.5-VL-72B 8.7 0.6 1.1 52.7 47.7 4.4 3.2 85.9
LLaVa-v1.6-34B 6.3 0.5 0.9 65.8 46.3 1.2 1.3 92.2
Claude 3.5 8.2 0.5 0.9 55.8 51.3 2.1 1.0 89.6
Visual-RAG Qwen2.5-VL-72B 6.5 0.5 1.0 74.0 42.3 2.7 5.2 11.0
LLaVa-v1.6-34B 5.3 0.4 0.8 76.8 44.1 3.2 4.5 12.0
Claude 3.5 6.0 0.6 1.1 85.1 39.5 1.8 6.1 14.5
WebQA Qwen2.5-VL-72B 8.8/8.1 0.8/16.1 1.5/10.8 62.8/70.4 52.5/7.9 6.5/3.6 5.9/13.5 89.7/30.4
LLaVa-v1.6-34B 6.5/7.1 0.7/15.8 1.3/9.8 64.5/70.8 50.2/6.2 2.5/4.1 1.1/12.2 91.4/28.6
Claude 3.5 9.0/7.9 1.1/1.8 1.7/3.7 26.6/98.8 62.7/10.7 0.3/0.6 72.0/8.8 87.4/13.6
Table 3: Diagnostic metrics on the Filtered split (image/text pairs where applicable).Best and Second-best are highlighted per row and split.All metrics are reported as percentage points except F1 (points on [0,100]).(Self-knowledge $\rightarrow$ Self-know.; Claim Recall $\rightarrow$ Claim R.; Context Precision $\rightarrow$ Ctx.Prec.)
Dataset Model $\DeltaCR Δ\DeltaCP$ VIS-Hit@k TXT-MissRate CMA V-HR D-HR
Chart-MRAG Phi-4 (GPD) 54.54 59.50 81.38 83.00 13.02 66.87 45.2
WebQA Phi-4 (GPD) 6.59 1.50 14.79 87.00 10.86 12.41 0.33
WebQA Claude 3.5 (GPD) 3.09 -0.35 13.00 83.33 6.52 32.44 3.70
[…]
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 InternVL3-8B Pixtral-12B-2409
DO GPD GO DO GPD GO DO GPD GO DO GPD GO DO GPD GO DO GPD GO
Chart-MRAG 22.7 25.4 26.8 10.3 18.1 13.9 8.2 15.5 9.7 15.9 19.8 17.4 23.3 19.5 21.7 14.8 14.1 16.6
MRAG-Bench 30.9 29.6 29.8 17.8 16.4 20.7 26.9 24.1 29.7 2.1 2.4 27.9 26.2 21.3 28.5 33.6 29.8 28.2
VisRAG-ArXiv 4.4 4.1 4.3 4.3 11.7 8.5 7.3 4.1 5.0 1.7 4.4 5.5 6.8 5.1 5.2 3.2 1.4 1.8
VisRAG-Doc 1.2 29.1 31.3 12.9 20.5 20.2 2.4 26.6 25.9 2.3 21.6 28.3 4.7 38.4 38.1 0.6 26.5 27.1
VisRAG-Plot 5.9 33.8 41.4 8.4 12.9 12.6 12.7 25.7 29.3 6.2 6.7 32.1 4.9 28.3 28.4 0.8 51.2 52.0
VisRAG-Slide 1.6 18.7 18.1 1.9 8.4 8.2 4.3 18.9 18.8 1.2 24.6 15.3 0.9 17.5 15.7 0.6 15.3 15.6
Visual-RAG 14.8 26.9 27.3 13.6 18.7 18.1 15.5 25.4 26.5 15.1 21.4 20.2 17.6 23.1 23.2 16.5 23.6 18.9
WebQA 17.2 15.8 15.4 12.3 12.6 13.4 9.9 14.3 16.4 17.5 7.3 12.8 20.1 20.8 17.4 25.6 25.3 0.9
Table 5: Scores (EM preferred; Accuracy used when EM is unavailable) under the three retrieval modes: DO = distractors_only, GPD = gt_plus_distractors, GO = gt_only. † marks the best in GO; ‡ marks the best in GPD; * marks the smallest $|\mathrm{GO}-\mathrm{GPD}|$ (most robust to distractors).
4.3 Retrieval Modes and Distractors
We compare distractors_only (DO), gt_plus_distractors (GPD), and gt_only (GO). Table 5 shows that DO degrades performance substantially; GPD lies between DO and GO. The GPD–GO gap quantifies susceptibility to distractors.
5 Experiments and Results
We evaluate MRAG-Suite along three axes: (1) end-task accuracy on short answers, (2) factuality and grounding quality of long answers via MM-RAGChecker, and (3) robustness to query difficulty, ambiguity, and distractors.
5.1 Experimental Setup
Models.
Our main baseline is LLaVa-v1.6-34B (vision-language model fine-tuned from LLaMA-2). Retrieval uses a CLIP-based dual encoder for images and a DPR retriever for text. At test time we fetch up to 5 passages and 5 images per query (when available) and serialize them (image features/captions + text) before generation. The model is prompted to output a short answer followed by a long answer. For comparison we also run Claude 3.5 (API) on a subset to estimate the upper bound of current proprietary VLMs.
Metrics.
For short answers we report EM/Accuracy. For long answers we compute ROUGE-L against our reference explanations and all claim-level diagnostics using MM-RAGChecker (Hallucination Rate, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, Modality Bias). We further report the fraction of retrieved items actually used (evidence utilization). Retrieval robustness is probed under three modes: gt_only (GO), gt_plus_distractors (GPD), and distractors_only (DO). Unless noted, we report macro averages on the Filtered split with decoding temperature =0=0 and identical retriever outputs across models.
5.2 End-task Accuracy (short-answer EM/Acc)
Table 2 reports EM/Accuracy across datasets and models. Overall accuracies are modest and vary widely by domain. On the Filtered split, the highest per-dataset scores appear on VisRAG-ArXiv (e.g., up to 63.9 with InternVL3-8B) and VisRAG-Plot (e.g., 51.4 with Pixtral-12B-2409), whereas Visual-RAG and VisRAG-Slide are the most challenging (typically in the 8–24 range across models). WebQA sits in the middle (best around 26.4 on the Filtered split), while Chart-MRAG remains difficult (best Filtered around 22.4).
Comparing Filtered vs. Full, the Full split is generally easier (e.g., Chart-MRAG improves substantially for several models, such as Pixtral-12B-2409 from 14.7 to 29.4), though the effect is not uniform across all datasets and models (e.g., VisRAG-ArXiv sometimes drops for stronger models when moving to Full). These trends confirm that our filtering increases non-triviality and retrieval dependence without relying on a single macro-average.
5.3 Claim-level Diagnostics (MM-RAGChecker)
Table 3 summarizes claim-level diagnostics on the Filtered split. We observe substantial variation across domains and models rather than a single global rate. Hallucination spans a wide range: it is comparatively low on document-like or chart-heavy sets (e.g., VisRAG-ArXiv around 30–41 for major models; Chart-MRAG can be as low as 24.9 on images for Claude 3.5) but very high on fine-grained visual recognition (Visual-RAG often 74–85). Correspondingly, Faithfulness is highest on structured pages/plots and lowest on open-world visual recognition.
Evidence utilization and cross-modality signals further explain these gaps. On WebQA, Context Precision for images is high (e.g., 89–91), while text exhibits much lower precision (often 28–31), indicating that retrieved passages are frequently unused or unhelpful. Across datasets we also see modality skew: models tend to commit to either text or images instead of reconciling both, which aligns with the elevated hallucination on visually challenging sets and the reduced per-claim F1 on those splits.
Overall, MM-RAGChecker reveals that (i) hallucinations remain frequent on visually demanding queries even when retrieval succeeds, (ii) many retrieved items are not credited as supporting any claim (low context precision, especially on text), and (iii) balancing visual and textual cues is a key bottleneck for long-form grounding.
5.4 Effect of Retrieval Modes
[…]
5.5 Prompt Sensitivity Study
We sweep 12 prompt designs (init style, few-shot exemplar count, reasoning pattern). Table 6 lists the best-performing prompt per dataset; Figure 3 visualizes the direct vs. retrieve-then-reason trend. Results in Table 6 are averaged over six models on the Filtered split with fixed seeds and decoding settings. Table 6 reports, for each dataset, the prompt that achieves the highest mean Accuracy/EM across six models (filtered split). We observe clear patterns: Reasoning matters: retrieve_then_reason dominates on text-heavy retrieval tasks (MRAG, WebQA), whereas direct is consistently better on VisRAG-* visual QA splits where context is already focused. Few-shot count is task-dependent: Using three exemplars (ex3) suffices—and sometimes outperforms ex6—for purely visual/numeric QA (plots/slides). In contrast, ex6 helps more on document-style and open-web questions that benefit from pattern coverage. Init style is secondary: Switching between plain and expert seldom changes the rank unless combined with the right reasoning pattern and example count.
Dataset Best Prompt Mean Acc. (%)
MRAG-Bench expert3 retrieve & reason 31.5
VisRAG-ArXiv expert6 direct 5.0
VisRAG-Doc plain6 direct 34.3
VisRAG-Plot expert3 direct 36.5
VisRAG-Slide plain3 direct 35.2
WebQA expert6 retrieve&reason 23.0
Table 6: Best-performing prompt (by mean EM/Acc across models) per dataset on the filtered split. The patterns highlight when “retrieve-then-reason” helps (text-heavy tasks) versus when direct prompting is sufficient (focused visual tasks).
Figure 3:
5.6 Impact of Distractors on Diagnostic Metrics
[…]
5.7 Effect of Filtering
Filtering not only reduces accuracy but also changes answer style. Filtering increases the average long-answer length from 45 to 60 tokens, yet factuality does not improve: MM-RAGChecker flags a larger fraction of clauses as unsupported on the Filtered split (cf. Table 3). This suggests that harder questions trigger verbosity rather than better grounding.
6 Conclusion
We presented MRAG-Suite, a comprehensive suite for evaluating and diagnosing visual retrieval-augmented generation in multimodal question answering. MRAG-Suite introduces a unified benchmark that stresses truly multimodal, knowledge-intensive QA, combining diverse sources like web images, textual articles, and informational charts. To focus evaluation on the intended challenges, we devised a two-step filtering method to remove trivial cases and emphasize multi-hop, vision+language reasoning. We also contributed an automated way to generate long-form answers with evidence, enabling novel evaluation of explanations in addition to factual answers. Crucially, we developed MM-RAGChecker, a claim-level multimodal verifier, to assess the factual consistency of generated answers.
Limitations
This work targets visual RAG for images, charts, slides, scanned pages, and scholarly PDFs; results may not transfer to videos, embodied settings, or low-resource languages. Hybrid retrieval depends on CLIP-like encoders and OCR text, so known dataset/model biases and OCR noise can affect measured recall/precision and inflate hallucination labels. Our diagnostics rely on LLM/VLM components for claim extraction and entailment; despite prompt controls and spot-checks, such judges can diverge from expert assessments in domain-specific cases. Scores are also sensitive to prompt format, retrieval $k$, and distractor sampling; we report settings but do not exhaust all combinations. Some baselines use third-party APIs, so provider updates and rate limits may introduce temporal variance despite fixed prompts and released scripts.
Appendix A Ambiguity Process
A.1 Ambiguity Definition and Annotation
We curate a 200-question subset to study ambiguity in multimodal RAG.
Definition.
A question is AMBIGUOUS if a reasonable reader cannot determine a single correct answer without additional constraints (time, location, unit, object, visibility in the image, etc.); otherwise it is CLEAR. For image-conditioned queries, we additionally mark as ambiguous when key visual evidence is not visible or is blurred.
Detection Pipeline.
We adopt a two-stage procedure:
  1. 1.
    LLM-based pre-filter. We prompt an assessor model to output CLEAR or AMBIGUOUS together with a short rationale (30 words). We use separate prompts for text-only and image-conditioned questions (see Appendix A.3).
  2. 2.
    Human adjudication. Two annotators independently review the flagged items and resolve disagreements. We achieve substantial agreement (Cohen’s $\kappa=0.74$).
Rewriting and Answer Normalization.
For each ambiguous item we:
  • Evidence-grounded answer rewrite: given retrieved evidence and the original answer, an LLM produces a concise answer or returns “Evidence inconclusive.” (Appendix A.5).
  • Query disambiguation: another prompt rewrites the question into a precise version that can be answered solely from the evidence (Appendix A.6).
The final ambiguity split (200 queries) and a matched clear set (200 queries) are used in Appendix. A.7 to isolate the impact of ambiguity on accuracy, claim recall, and hallucination rate.
A.2 Error Patterns under Ambiguity
On the 200-query ambiguity subset, two dominant patterns emerge:
  1. 1.
    Single-interpretation bias: the model commits to one reasonable sense, lowering Claim Recall and short-answer accuracy.
  2. 2.
    Interpretation conflation: retrieval returns evidence for multiple senses; the generator merges them, producing composite neutral claims and higher hallucination rates.
These results motivate an ambiguity detection & clarification step before final answer generation.
A.3 Ambiguity Detection (Image-conditioned)
A.4 Ambiguity Detection (Text-only)
A.5 Evidence-based Answer Rewrite
A.6 Ambiguous Query Rewrite
Table 7: Ambiguity subset construction statistics. We report candidate counts, items flagged as ambiguous by the two-stage pipeline, and the final curated set (200).
Candidates Marked Ambiguous Final Set
Text-only 26,120 13,823 184
Image-conditioned 4,557 1,224 16
Total 30,677 15,047 200
A.7 Ambiguity Subset Results
On the 200-query ambiguous subset, short-answer EM drops by 5.9 points compared to clear queries, while hallucination rate increases by 9.5 points. Claim Recall and Context Precision also decrease, indicating single-interpretation bias or interpretation conflation; see Appendix. A.2 for qualitative patterns. Table 8 reports the aggregate metrics. Also, we have the ambiguous set in the code provided.
Table 8: Results on the 200-query ambiguity subset versus a matched clear set. We report EM/Accuracy, Hallucination Rate (lower is better), Claim Recall, and Context Precision, isolating the effect of query underspecification on retrieval and grounding.
EM/Acc $\uparrow$ Halluc. $\downarrow$ Claim Recall $\uparrow$ Ctx. Prec. $\uparrow$
Clear (200 random) 23.5 49.2 12.4 14.1
Ambiguous (200) 17.6 58.7 9.1 12.5
$\Delta$ (Ambig - Clear) -5.9 +9.5 -3.3 -1.6
Appendix B Retrieve system settings
Dataset Text Retriever Image Retriever
WebQA SBERT (mpnet) CLIP ViT-B/32
Chart-RAG DPR (ctx-enc) CLIP ViT-B/32
MRAG CLIP ViT-B/32 (img+txt avg)
VisRAG VisRAG-Ret
VisualRAG CLIP ViT-B/32
Table 9: Backbones used in the hybrid retriever for each sub-corpus. Images are indexed with CLIP variants; text uses DPR/SBERT; VisRAG documents use a document-image retriever.
Appendix C 12 Prompt Experiment Settings
Layer Arg name Values in this paper Template / Semantics
System Prompt init_style plain plain: “Please read the following question and retrieve the relevant document(s)/image(s) to produce the best possible answer.”
expert expert: “You are an expert in <DOMAIN>. Analyze the question carefully, think step-by-step about which sources to use, and present your answer clearly at the end.”
Few-shot Demo example_style ex1, ex3, ex6 Three fixed 0-shot style exemplars (chosen from the pool). Content differs in wording/format hints.
max_examples 0 or 5 Whether to insert 5 illustrative QA pairs (kept at 0 in our sweep).
Reasoning Instruction direct direct: “Question: … Answer:” (no mandated reasoning steps).
reasoning_process retrieve_then_reason: “First outline retrieval/reasoning, then give final answer.”
Answer Format answer_format free Free-form text (no forced JSON or letter).
Context Packaging context_order img_first Images precede text in the prompt (fixed in our runs).
include_doc_ids 0 / 1 Whether to show doc/image IDs (0 in our runs).
Table 10: Prompt dimensions and concrete templates used in our 12-way sweep.
Nickname init_style ex_style Example template
expert1 expert ex1 Q: {q}
A: {ans}
expert2 expert ex3 Question + brief Explanation + Final Answer
expert3 expert ex6 Dialog style: User: {q} / Assistant: {ans}
plain1 plain ex1 Q: {q}
A: {ans}
plain2 plain ex3 Question + brief Explanation + Final Answer
plain3 plain ex6 Dialog style: User: {q} / Assistant: {ans}
Table 11: Mapping from prompt nicknames to template families used in the 12-way prompt sweep. Variations differ in initialization style and exemplar formatting.
p_reason Reasoning block (excerpt) COT line?
direct Answer directly without explaining your reasoning. No
retrieve_then_reason First list which retrieved snippets/images you will use (by Doc/Image IDs). Then reason step by step using them, and finally give the answer. Yes
structured Follow this structure: 1) Question understanding 2) Relevant sources (IDs or short quotes) 3) Reasoning 4) Final Answer Yes
plan_execute Outline a short plan first (bullet points). After the plan, execute it and derive the final answer. Yes
verify Propose an initial answer. Then self-check it briefly. If it passes, output the final answer prefixed with “FINAL:”. Yes
none (no additional block) No
Table 12: Reasoning patterns injected into prompts. Variants differ in whether to expose chain-of-thought–like structure and explicit self-verification before the final answer.
Dataset Qwen2.5-VL-72B LLaVa-v1.6-34B Claude 3.5 Phi-4 Pixtral-12B-2409 InternVL3-8B
Filt Full Filt Full Filt Full Filt Full Filt Full Filt Full
Chart-MRAG 22 19 15 25 10 11 13 12 14 29 19 27
MRAG-Bench 32 33 19 24 28 26 23 24 38 37 21 21
VisRAG-ArXiv 55 48 34 34 35 35 26 27 43 43 63 55
VisRAG-Doc 31 35 20 21 28 29 24 22 26 26 39 34
VisRAG-Plot 36 45 12 13 35 36 6 9 51 56 33 36
VisRAG-Slide 18 25 8 10 17 22 24 22 18 15 20 22
Visual-RAG 13 13 9 11 11 10 8 9 12 13 11 12
WebQA 16 14 12 13 14 17 10 11 26 27 17 28
Table 13: EM/Accuracy performance (%) of each model on the Filtered and Full splits.All numbers are percentage points; higher is better.
Appendix D Distractor results overview
Dataset Model Distractor Recall Precision F1 Halluc. Faith. Self-know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
VisRAG-ArXiv Qwen2.5-VL-72B distractors only 21.8 6.7 7.9 66.2 23.5 5.9 20.4 31.0
gt_plus_distractors 29.6 8.9 10.8 60.1 28.9 7.1 24.1 33.5
gt_only 33.4 10.8 12.7 55.3 32.1 8.2 25.9 35.8
LLaVa-v1.6-34B distractors only 18.4 5.9 6.8 69.4 20.1 4.8 18.0 28.2
gt_plus_distractors 24.7 7.6 9.1 63.5 24.9 5.9 21.6 30.4
gt_only 28.9 9.2 10.9 58.2 29.6 6.8 23.7 33.1
Claude 3.5 distractors only 30.4 9.2 10.1 61.4 28.0 7.6 29.9 40.0
gt_plus_distractors 40.7 10.6 14.3 56.0 37.0 7.0 25.6 33.0
gt_only 35.9 8.6 11.3 57.7 36.5 4.7 27.6 39.0
Phi-4 distractors only 22.5 15.3 13.7 50.8 20.6 12.9 22.2 33.8
gt_plus_distractors 34.4 24.0 23.5 45.3 22.3 16.3 26.0 36.0
gt_only 37.8 27.2 26.6 41.1 24.7 18.5 27.9 37.9
Pixtral-12B-2409 distractors only 17.6 7.1 6.3 70.5 18.7 5.1 16.0 27.6
gt_plus_distractors 25.9 10.4 9.7 63.0 23.3 7.4 18.9 29.5
gt_only 30.5 12.8 12.0 58.9 25.6 8.3 21.5 31.2
InternVL3-8B distractors only 26.5 14.6 12.6 59.3 24.0 10.8 28.8 37.0
gt_plus_distractors 39.2 15.7 16.5 54.4 30.0 10.6 29.0 38.0
gt_only 40.3 14.6 17.6 48.2 27.4 9.5 24.4 35.0
VisRAG-Doc Qwen2.5-VL-72B distractors only 10.8 4.0 3.6 57.9 13.4 2.6 21.0 23.5
gt_plus_distractors 46.7 31.9 27.2 28.1 22.7 20.4 25.9 27.0
gt_only 52.9 36.2 31.0 24.4 24.5 22.6 24.1 25.1
LLaVa-v1.6-34B distractors only 9.4 3.7 3.2 60.1 12.1 2.3 19.5 22.0
gt_plus_distractors 39.8 26.1 22.0 33.9 20.1 16.4 23.8 25.6
gt_only 44.0 28.9 24.7 30.2 22.4 18.5 22.9 24.2
Claude 3.5 distractors only 12.0 4.6 4.2 38.2 16.9 3.0 24.5 26.0
gt_plus_distractors 75.5 35.2 38.2 41.4 32.2 25.4 27.7 31.0
gt_only 68.8 33.6 33.7 41.5 30.7 23.8 26.5 28.0
Phi-4 distractors only 18.1 12.2 11.0 27.6 7.5 10.5 20.8 22.7
gt_plus_distractors 45.3 37.8 37.0 15.2 4.7 32.3 24.5 26.0
gt_only 50.6 41.9 41.1 13.0 3.9 35.2 23.6 24.8
Pixtral-12B-2409 distractors only 7.8 3.5 3.0 61.3 10.6 2.1 18.8 21.3
gt_plus_distractors 34.4 25.6 20.4 36.8 17.7 15.1 22.1 23.9
gt_only 39.1 29.7 24.1 32.6 19.8 17.2 21.5 22.8
InternVL3-8B distractors only 8.0 4.1 3.4 55.2 12.8 3.1 22.5 25.0
gt_plus_distractors 68.3 65.2 26.5 15.8 25.6 45.5 28.0 26.5
gt_only 60.3 77.8 57.8 17.2 27.3 54.5 21.5 23.0
MRAG-Bench Qwen2.5-VL-72B distractors only 22.6 6.2 5.1 79.4 14.3 4.0 7.3 4.5
gt_plus_distractors 28.4 7.9 6.7 75.8 16.2 4.7 8.9 5.2
gt_only 33.7 9.5 8.4 71.1 18.6 5.3 10.4 6.0
LLaVa-v1.6-34B distractors only 21.1 5.3 4.1 82.4 13.2 3.5 8.2 4.2
gt_plus_distractors 27.8 4.8 3.1 77.1 17.6 4.3 10.1 6.0
gt_only 31.9 6.7 4.6 73.4 19.1 4.8 9.8 5.0
Claude 3.5 distractors only 42.2 3.9 4.6 73.5 19.8 3.8 9.2 5.5
gt_plus_distractors 47.2 2.8 3.9 80.5 16.1 2.4 7.7 6.0
gt_only 47.7 4.2 5.6 78.8 17.1 4.1 5.5 3.0
Phi-4 distractors only 14.5 12.6 10.8 36.4 8.8 11.9 7.1 4.6
gt_plus_distractors 18.7 18.3 16.2 31.4 9.3 16.3 8.6 5.5
gt_only 22.6 21.4 18.7 27.8 10.4 18.5 9.5 6.2
Pixtral-12B-2409 distractors only 12.8 6.9 5.7 84.6 11.0 5.6 6.4 3.9
gt_plus_distractors 17.2 8.8 7.2 80.3 12.5 6.4 7.3 4.2
gt_only 21.0 10.6 8.6 76.1 14.3 7.2 8.4 4.9
InternVL3-8B distractors only 35.3 37.6 31.0 54.5 9.3 35.1 5.3 3.0
gt_plus_distractors 29.0 33.1 25.1 46.7 8.2 32.1 5.3 3.5
gt_only 31.8 31.6 25.6 62.0 7.5 28.6 6.2 4.0
VisRAG-Plot Qwen2.5-VL-72B distractors only 10.1 3.9 2.6 61.7 12.2 2.7 9.4 10.2
gt_plus_distractors 17.6 7.2 5.8 58.5 11.6 5.4 9.7 9.6
gt_only 23.8 10.7 8.2 56.0 12.9 7.7 8.6 8.8
LLaVa-v1.6-34B distractors only 7.9 3.1 1.9 64.9 10.7 2.1 9.1 10.4
gt_plus_distractors 13.4 6.1 4.0 61.2 11.5 3.9 9.5 9.8
gt_only 19.6 8.3 5.7 58.9 12.1 5.5 8.9 9.3
Claude 3.5 distractors only 8.0 4.1 2.1 55.1 10.1 2.8 10.5 11.0
gt_plus_distractors 19.8 9.4 6.6 77.8 13.4 8.9 9.0 9.0
gt_only 26.5 11.6 9.4 75.2 12.7 11.1 7.5 8.0
Phi-4 distractors only 11.3 4.7 1.8 83.7 10.3 4.0 10.0 10.0
gt_plus_distractors 16.8 8.9 3.7 79.5 11.2 6.8 9.6 9.7
gt_only 22.4 11.5 5.2 76.4 12.0 9.4 8.4 8.7
Pixtral-12B-2409 distractors only 6.5 2.8 1.1 66.0 9.8 1.7 8.8 9.5
gt_plus_distractors 12.9 6.4 2.9 62.1 10.5 3.7 9.2 9.6
gt_only 18.7 8.9 4.6 59.6 11.2 5.2 8.7 9.0
InternVL3-8B distractors only 2.3 1.8 0.4 52.3 11.0 1.7 8.5 9.0
gt_plus_distractors 25.5 15.1 7.9 72.9 9.3 13.8 10.0 10.0
gt_only 21.5 11.2 3.2 76.5 8.2 10.3 11.1 12.0
VisRAG-Slide Qwen2.5-VL-72B distractors only 12.6 5.5 4.8 63.3 14.7 3.8 51.1 56.8
gt_plus_distractors 41.3 27.2 26.1 36.5 42.2 21.1 46.9 50.6
gt_only 49.7 34.9 33.5 31.2 48.4 23.2 43.8 48.0
LLaVa-v1.6-34B distractors only 10.9 4.9 4.1 66.0 13.3 3.3 49.6 54.3
gt_plus_distractors 35.6 22.7 21.0 40.2 39.1 18.8 45.1 49.2
gt_only 42.8 29.6 27.4 35.5 43.7 20.6 41.9 46.6
Claude 3.5 distractors only 23.1 12.0 11.2 53.3 36.7 9.1 53.3 59.1
gt_plus_distractors 78.8 30.6 37.3 38.7 45.8 15.4 55.7 61.1
gt_only 64.2 39.1 36.0 30.1 49.2 18.9 49.2 56.0
Phi-4 distractors only 19.9 11.7 10.3 41.2 31.6 8.6 50.1 55.6
gt_plus_distractors 71.8 42.7 44.5 33.7 40.1 23.4 45.9 48.6
gt_only 60.4 47.9 45.0 29.0 37.7 26.1 43.6 46.2
Pixtral-12B-2409 distractors only 9.6 5.1 4.2 67.9 12.0 3.4 48.5 53.7
gt_plus_distractors 29.8 18.9 16.7 45.6 34.7 16.3 44.3 47.5
gt_only 36.7 24.5 22.1 39.8 38.9 17.9 41.1 45.3
InternVL3-8B distractors only 17.2 4.4 5.1 58.7 34.0 3.2 54.1 59.7
gt_plus_distractors 67.1 42.6 41.9 23.4 51.4 22.5 47.3 51.4
gt_only 57.8 41.1 38.7 28.8 44.6 23.8 40.8 45.8
WebQA Qwen2.5-VL-72B distractors only 14.0/13.1 6.5/5.8 5.0/4.3 69.1/71.2 9.4/8.6 5.0/5.4 12.8/12.0 14.0/13.2
gt_plus_distractors 18.3/17.2 8.9/8.1 7.3/6.5 66.5/68.7 11.2/10.3 6.3/6.8 12.0/11.3 13.5/12.6
gt_only 22.4/21.0 11.3/10.2 9.1/8.0 62.8/64.9 13.0/12.0 7.4/8.0 11.2/10.5 12.7/11.8
LLaVa-v1.6-34B distractors only 12.6/11.4 5.7/5.0 4.5/3.8 71.6/73.9 8.0/7.2 4.8/5.1 12.1/11.3 13.1/12.1
gt_plus_distractors 16.8/15.6 8.2/7.4 6.6/5.7 68.7/70.8 10.4/9.5 5.9/6.3 11.3/10.6 12.7/11.8
gt_only 21.3/19.9 10.5/9.4 8.4/7.3 65.2/67.3 12.6/11.6 7.0/7.6 10.7/10.0 12.0/11.2
Claude 3.5 distractors only 23.0/21.5 4.7/4.3 5.2/4.4 63.4/86.4 31.9/4.1 3.7/3.5 12.6/4.8 15.0/12.0
gt_plus_distractors 22.4/20.7 5.9/4.3 5.4/4.7 61.1/87.2 30.9/4.4 4.9/3.3 10.1/7.0 13.0/16.5
gt_only 26.8/24.1 6.8/5.0 6.6/5.4 57.6/84.3 33.5/5.2 5.5/3.9 11.2/8.1 14.2/17.8
Phi-4 distractors only 14.2/12.7 22.0/19.1 8.7/7.5 53.4/60.2 9.1/0.7 20.1/18.5 10.9/4.7 13.6/12.3
gt_plus_distractors 10.4/12.5 25.6/22.3 9.5/11.1 49.7/56.5 8.4/0.5 23.9/22.0 11.8/5.2 14.5/13.0
gt_only 16.1/14.8 29.0/24.7 12.1/12.4 46.2/53.0 10.2/0.9 26.7/24.1 12.3/5.8 15.9/13.9
Pixtral-12B-2409 distractors only 10.5/ 9.6 15.3/14.1 6.4/5.7 58.9/65.2 6.9/1.0 13.8/12.6 9.6/3.9 12.0/10.9
gt_plus_distractors 13.4/12.6 19.6/17.4 8.7/7.5 55.2/61.8 8.0/1.2 17.3/15.5 10.5/4.4 13.2/11.9
gt_only 17.4/16.0 23.5/20.7 11.3/9.8 51.7/58.1 9.6/1.6 20.4/18.0 11.8/5.0 14.7/13.0
InternVL3-8B distractors only 27.9/27.3 13.8/14.6 14.7/12.8 68.1/77.2 13.8/8.8 13.1/11.9 13.0/4.5 14.0/10.5
gt_plus_distractors 32.0/31.0 15.4/15.3 17.0/15.1 69.0/74.4 14.2/8.4 14.8/12.2 11.2/6.3 14.0/14.0
gt_only 16.3/16.9 7.0/6.9 5.6/5.1 64.0/63.6 8.9/12.0 6.1/5.3 8.5/6.4 11.5/13.5
Table 14: Full $M$-RAGChecker diagnostics across datasets, models, and retrieval modes. For each setting we report per-modality (image/text) Recall, Precision, F1, Hallucination, Faithfulness, Self-Knowledge, Claim Recall, and Context Precision.
Dataset Model Distractor Recall Precision F1 Halluc. Faith. Self‑know. Claim R. Ctx.Prec.
(img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt) (img/txt)
ChartMRAG-Bench Qwen2.5-VL-72B distractors only 11.0/10.0 4.2/3.7 2.9/2.4 60.6/62.8 12.8/11.9 2.6/2.9 9.8/9.1 10.6/10.0
gt_plus_distractors 18.4/17.1 7.6/6.9 6.2/5.4 57.1/59.4 12.2/11.1 5.1/5.6 10.1/9.4 10.0/9.7
gt_only 24.6/23.1 11.1/10.3 8.6/7.7 54.6/56.9 13.4/12.6 7.4/8.1 9.0/8.3 9.2/8.9
LLaVa-v1.6-34B distractors only 8.6/7.6 3.4/2.9 2.2/1.7 63.7/65.8 11.2/10.5 2.0/2.3 9.5/8.8 10.7/10.2
gt_plus_distractors 14.2/13.0 6.5/5.8 4.3/3.5 60.1/62.3 12.0/11.0 3.8/4.2 9.9/9.2 10.2/9.8
gt_only 20.7/19.2 8.9/8.1 6.1/5.3 57.8/60.1 12.6/11.8 5.4/6.0 9.2/8.6 9.6/9.1
Claude 3.5 distractors only 8.6/7.5 4.3/3.9 2.3/1.9 54.0/56.0 10.4/9.8 2.7/3.0 10.8/10.2 11.2/10.7
gt_plus_distractors 20.7/18.9 9.8/8.9 6.9/6.2 76.6/78.9 13.8/12.7 9.2/9.8 9.2/8.9 9.2/8.8
gt_only 27.5/25.4 12.2/11.0 9.7/8.9 74.3/76.6 13.1/12.2 11.2/11.9 7.7/7.3 8.2/7.8
Phi-4 distractors only 12.1/10.6 4.9/4.4 2.0/1.7 82.4/84.7 10.9/9.8 3.7/4.0 10.2/9.7 10.2/9.9
gt_plus_distractors 17.6/16.0 9.4/8.5 4.1/3.4 78.3/80.6 11.7/10.8 6.5/7.0 9.9/9.3 9.9/9.5
gt_only 23.3/21.6 12.0/10.9 5.5/4.9 75.1/77.4 12.5/11.6 9.1/9.7 8.7/8.1 8.9/8.5
Pixtral-12B-2409 distractors only 7.2/6.0 3.1/2.6 1.3/1.0 64.8/66.9 10.3/9.4 1.6/1.9 9.1/8.5 9.8/9.3
gt_plus_distractors 13.7/12.2 6.7/5.9 3.1/2.7 60.9/63.1 11.0/10.1 3.6/3.9 9.5/8.9 9.9/9.4
gt_only 19.6/18.0 9.3/8.2 4.9/4.3 58.6/60.8 11.7/10.8 5.1/5.6 9.1/8.5 9.3/8.9
InternVL3-8B distractors only 2.6/2.1 1.9/1.6 0.5/0.4 51.1/53.5 11.3/10.7 1.6/1.8 8.7/8.3 9.2/8.8
gt_plus_distractors 26.5/24.6 15.7/14.4 8.4/7.5 71.8/74.0 9.7/8.9 14.3/15.0 10.1/9.7 10.1/9.8
gt_only 22.3/20.8 11.6/10.7 3.5/2.9 75.4/77.5 8.5/7.8 10.6/11.1 11.4/10.8 12.1/11.4
Visual-RAG Qwen2.5‑VL‑72B distractors only 18.4 6.3 6.8 42.1 27.2 48.3 4.9 18.2
gt_plus_distractors 44.5 14.7 16.8 36.3 34.5 53.9 6.8 25.0
gt_only 41.3 13.2 15.4 37.5 32.8 51.1 6.5 23.3
LLaVa‑v1.6‑34B distractors only 16.1 5.2 5.7 43.7 25.4 45.0 4.4 16.1
gt_plus_distractors 42.0 12.9 14.7 38.8 30.1 51.6 6.0 22.7
gt_only 38.8 11.8 13.2 39.5 28.0 50.2 5.7 21.4
Claude 3.5 distractors only 23.3 7.0 8.4 40.2 28.6 50.3 6.2 20.0
gt_plus_distractors 52.7 16.5 19.7 35.0 36.7 55.5 7.5 28.8
gt_only 48.1 15.2 17.4 36.1 34.9 53.2 7.1 26.5
Phi‑4 distractors only 20.7 6.5 7.3 39.0 26.0 46.4 5.3 18.0
gt_plus_distractors 47.2 15.0 17.6 34.1 32.5 52.9 6.7 25.2
gt_only 43.6 13.9 16.0 35.3 31.0 50.7 6.4 23.7
Pixtral‑12B‑2409 distractors only 15.3 5.8 6.2 44.9 23.5 43.2 4.1 15.3
gt_plus_distractors 40.8 12.0 13.9 38.2 29.8 49.8 5.8 21.9
gt_only 37.9 11.1 12.8 39.0 27.7 48.1 5.5 20.4
InternVL3‑8B distractors only 22.6 8.1 9.2 41.3 30.0 49.0 6.0 19.5
gt_plus_distractors 50.1 17.2 20.3 33.7 37.8 54.4 7.8 27.1
gt_only 45.7 15.5 18.1 34.9 35.3 52.0 7.3 25.0
Table 15: